### v1

# NLP for Social Good: A Survey of Challenges, Opportunities, and Responsible Deployment

Abstract
Recent advancements in large language models (LLMs) have unlocked unprecedented possibilities across a range of applications. However, as a community, we believe that the field of Natural Language Processing (NLP) has a growing need to approach deployment with greater intentionality and responsibility. In alignment with the broader vision of AI for Social Good Tomašev et al. (2020), this paper examines the role of NLP in addressing pressing societal challenges. Through a cross-disciplinary analysis of social goals and emerging risks, we highlight promising research directions and outline challenges that must be addressed to ensure responsible and equitable progress in NLP4SG research.
NLP for Social Good: A Survey of
Challenges, Opportunities, and Responsible Deployment
Antonia Karamolegkou1, Angana Borah2, Eunjung Cho3, Sagnik Ray Choudhury4, Martina Galletti5,6, Rajarshi Ghosh7, Ishani Mondal8, Pranav Gupta9, Oana Ignat10, Priyanka Kargupta11, Neema Kotonya12, Hemank Lamba12, Sun-Joo Lee13, Arushi Mangla10, Deniz Nazarova14, Poli Nemkova4,13, Dina Pisarevskaya15, Naquee Rizwan16, Nazanin Sabri17, Dominik Stammbach18, Anna Steinberg19,20, David Tomás21, Steven R Wilson22, Bowen Yi2,23, Jessica Zhu8, Arkaitz Zubiaga15, Anders Søgaard1, Alexander Fraser20,27, Zhijing Jin3,24,25,26, Rada Mihalcea2, Joel R. Tetreault12, Daryna Dementieva27 1University of Copenhagen 2University of Michigan-Ann Arbor 3ETH Zurich 4 University of North Texas 5Sony Computer Science Laboratories - Paris 6University of Rome “La Sapienza” 7Algoverse 8University of Maryland, College Park 9Lowe’s 10Santa Clara University 11University of Illinois Urbana-Champaign 12Dataminr 13 United Nations Development Programme (UNDP) 14University of Washington 15Queen Mary University of London 16IIT Kharagpur 17University of California San Diego 18 Princeton University 19LMU Munich 20Munich Center for Machine Learning (MCML) 21University of Alicante 22University of Michigan-Flint 23University of Southern California 24Max Planck Institute for Intelligent Systems, Tübingen 25Vector Institute 26University of Toronto 27Technical University of Munich Correspondence: antka@di.ku.dk, daryna.dementieva@tum.de
1 Introduction
“Understanding the problem is half the solution.”
— Charles Kettering
Figure 1: Mapping NLP applications for Social Good (NLP4SG) with global goals and risks.
To fully realize the potential of NLP, it is essential to look beyond technical achievements and reframe tasks around pressing societal needs. We draw on insights from the United Nations Sustainable Development Goals [^1] https://sdgs.un.org/goals (UN SDGs) and the 2025 Global Economic Risks Report [^2] https://www.weforum.org/publications/global-risks-report-2025/digest (GR) to provide a foundation for an interdisciplinary recontextualization of NLP, encouraging reflection on how language technologies intersect with today’s most pressing challenges. We selected these two agendas as, from a social good perspective, UN SDGs offer a global framework for fostering peace and prosperity for people and the planet. However, while highly influential, these goals were established in 2015—prior to the rapid advancements in artificial intelligence. To contextualize them within today’s technological landscape, we also draw on insights from the 2025 GR Report, which highlights both the transformative potential and the emerging global risks (GRs) associated with technology and information processing. Our effort builds on prior research that assesses the role of NLP through positive impact Hovy and Spruit (2016); Jin et al. (2021), maps NLP4SG work to the SDGs Adauto et al. (2023); Gosselink et al. (2024), outlines open questions in modern NLP Ignat et al. (2024b), and limitations in NLP and AI pipelines Mihalcea et al. (2025). Thus, our research goal in this work is threefold:
  1. RQ1:
    What NLP-based solutions already support positive social impact?
  2. RQ2:
    What are the main challenges in developing NLP for positive social impact?
  3. RQ3:
    What promising directions are currently overlooked?
To address these questions, we identify nine key NLP research directions, as illustrated in Figure 1, using the relevant SDGs and GRs. The following sections summarize research tasks, [^3] Details about methodology and motivation of paper selection can be found in the Appendix A. the open challenges and opportunities within each domain, concluding with overarching open research directions and recommended community actions for proactive NLP4SG efforts.
2 [Uncaptioned image]  Healthcare
NLP in healthcare is a space that targets problems associated with SDG3: Good Health and Well-being by improving healthcare delivery and outcomes. It also targets critical global risks (GR4), including the decline in health and well-being, health workforce shortages, and infectious disease outbreaks.
Mental Health
Given that the need for mental health support increasingly remains unmet, NLP systems, specifically LLMs, address an access gap for users who are limited in time and resources Tong et al. (2023); Hua et al. (2024). As counselors, they could assist in detecting conditions like depression and addiction from clinical or social media data Giuntini et al. (2020); Yang et al. (2023a); responding empathetically by interpreting emotion and generating therapeutic dialogues Shen et al. (2020); Grandi et al. (2024); and tracking user mood or crises over time Čosić et al. (2024); Gong et al. (2019). As clients, NLP tools simulate diverse personas to train and evaluate counselors Louie et al. (2024); Liu et al. (2025a).
Physical Health
Prior work has focused on physical well being, using social media mining for tracking physical activity, sleep patterns Sakib et al. (2021); Shakeri Hossein Abad et al. (2022), diet habits van Erp et al. (2021); Hu et al. (2023), gauging public health attitudes such as mask-wearing during COVID-19 He et al. (2021), and detecting risky behaviors like substance use Hu et al. (2021); Lin et al. (2023). In clinical settings, NLP aids in medical record analysis via classification for treatment decisions, named entity recognition for patient-trial matching, relation extraction linking symptoms and treatments, predictive modeling of treatment responses (Jerfy et al., 2024), and information extraction to organize text reports Sheikhalishahi et al. (2019); Landolsi et al. (2023).
Challenges:
Key challenges include health-related data, which is often scarce, sensitive, and affected by systemic biases, with language limitations and underrepresentation of marginalized groups raising ethical and privacy concerns (Ford et al., 2019; Shakeri Hossein Abad et al., 2022; Gunal et al., 2025); evaluation, which must go beyond accuracy to reflect fairness, contextual understanding, human-centered values (e.g. empathy), and ensure reproducibility; and long-term user impact, as reliance on LLMs for sensitive tasks risks sycophancy, ELIZA effects Ekbia (2008), and overdependence. There is also a lack of causal frameworks and interpretable models, which complicates the understanding in the outcomes of NLP systems Zhang et al. (2022). Remote care with LLM-based tools risks diverting individuals from essential in-person treatment Khawaja and Bélisle-Pipon (2023); Sweeney et al. (2021). This highlights a broader challenge: ensuring such technologies support—not replace—human professionals, with responsible use as a guiding principleShakeri Hossein Abad et al. (2022); Brown and Halpern (2021).
Opportunities:
Future work should extend beyond text to multimodal approaches that integrate speech prosody, facial expressions, physiological signals, and other data sources like sensors, wearables, or images for richer context and personalization Puce et al. (2025). Multi-agent and adaptive dialogue systems that consider user history, emotions, and culture could boost performance. Designing holistic evaluation frameworks in real-world simulations to ensure privacy, explainability, fairness, and accessibility can help address AI risks Lawrence et al. (2024); Yao et al. (2024). At the policy level, NLP can analyze public sentiment on AI used for healthcare applications from actual users, and health guidelines to inform better regulations and health campaigns Lindquist et al. (2021). Crucially, interdisciplinary research is needed to build AI-augmented therapeutic frameworks that complement human care—especially for underserved communities.
3 [Uncaptioned image]  Education
The integration of LLMs into education offers tools that support SDG4: Quality Education, and also addresses a variety of risks by fostering informed and critical individuals (GR4, GR5). NLP is used for automated feedback Jurenka et al. (2024); Bauer et al. (2023); Gao et al. (2024); Stamper et al. (2024); Ramesh and Sanampudi (2022), tailored support Kazemitabaar et al. (2024); Daheim et al. (2024), and self-paced learning Kazemitabaar et al. (2023). NLP tools can expand access in underserved regions Yu et al. (2024), bridge language gaps Molina et al. (2024); Kwak and Pardos (2024), assist learners with disabilities Cheng et al. (2024), and ease teacher workload Lan and Chen (2024); Wang et al. (2024b); Shridhar et al. (2022).
AI Literacy
As AI systems become more embedded in everyday life, AI literacy is essential. A recent review (Yang et al., 2025) draws connections with digital, data, and algorithmic literacies. However, NLP-driven efforts for AI literacy remain scattered (Long and Magerko, 2020; Congress, 2023; Moorkens et al., 2024; Tapo et al., 2025; Korea Education and Research Information Service, 2025)
Challenges:
Model limitations—such as lack of pedagogical reasoning Wang and Demszky (2023); Macina et al. (2023, 2025), misaligned explainability Okolo and Lin (2024), and accuracy Stamper et al. (2024); Kargupta et al. (2024)—hinder the effective integration of NLP into education. Mixed perceptions and mistrust toward AI also remain a barrier Nader et al. (2022); Laupichler et al. (2024). Broader issues like language, cultural differences, curriculum gaps, weak policy support, and limited infrastructure—especially in developing regions—further restrict equitable access to AI/NLP in education Kathala and Palakurthi (2025).
Opportunities:
Further research should be encouraged for specific groups, such as teachers Du et al. (2024), students Shen and Cui (2024), and other professionals Lo (2024). Another important opportunity is the alignment with expert-annotated pedagogical traces and grounding evaluations in curriculum-linked outcomes, to bridge the gap between subject expertise and pedagogical effectiveness Macina et al. (2025); Lucy et al. (2024). Human-in-the-loop approaches have also shown promise in scaling expert strategies while preserving teacher agency, particularly benefiting underserved communities Wang et al. (2024c). Multi-agent simulations offer privacy-preserving environments to test classroom policies and assess equity metrics before real-world deployment Zhang et al. (2025). For example, Socratic planning agents encourage critical thinking over rote memorization, fostering deeper learning experiences Kargupta et al. (2024). Finally, community-driven initiatives are crucial for consolidating best practices for the development of safe and accountable educational AI systems Chu et al. (2025); Wen et al. (2024).
4 [Uncaptioned image]  Poverty
Economic downturn (GR3) is the sixth highest-ranked GRs, with serious implications for poverty—one of the world’s most pressing challenges (Lister, 2021) and the focus of SDG1. Nearly 700 million people continue to live on less than \$2.15 per day (Hasell et al., 2024; World Bank Group, 2024). NLP methods have been used to extract socioeconomic patterns from news (Lampos et al., 2014), analyze text (Paterson and Gregory, 2018; Hoeschle et al., 2025), classify poverty status Muñetón-Santa et al. (2022), extract poverty-related dimensions from interviews (Muñetón-Santa and Orozco-Arroyave, 2023) or analyze global narratives around poverty (Curto et al., 2024), and to identify performance disparities among different socioeconomic groups (Cercas Curry et al., 2024; Nwatu et al., 2023).
Challenges:
A major challenge is the fact that poverty data are often scarce or incomplete Tingzon et al. (2019); Fatehkia et al. (2020). Indicators like income or poverty status are rarely shared, making it hard to infer them from text. As a result, most socio-economic NLP studies rely on proxies such as mean income Hasanuzzaman et al. (2017); Abraham et al. (2020), education Cercas Curry et al. (2024), or (un)employment Preoţiuc-Pietro et al. (2015b, a). These proxies may not accurately reflect true poverty levels and often vary across studies, resulting in limited comparability and impact.
Opportunities:
To enable large-scale, global analysis of poverty, high-quality datasets labeled with income, socioeconomic status, or poverty indicators are essential, potentially gathered via data donations, user surveys, or social media statistics. [^4] https://datareportal.com/social-media-users Additionally, use-case-specific NLP applications can advance poverty research (Adauto et al., 2023). For example, model analysis can track the performance of current systems across socio-economic levels, information extraction can monitor government funding for poverty alleviation, and machine translation can improve resource access for non-English speakers. Finally, future work should develop clearer guidelines and taxonomies for categorizing poverty-related data and NLP methods to ensure consistency and comparability across studies.
5 [Uncaptioned image]  Peacebuilding
Peacebuilding is essential to achieving SDG16, and despite escalating threats like state-based armed conflict (R1); NLP’s application in key tools for peacebuilding like conflict prediction, human rights monitoring, and physical safety remains limited and underexplored.
Human Rights Violations
NLP in human rights mainly involves classification and detection of violations across languages, such as Arabic Twitter (Alhelbawy et al., 2016), Jordanian Arabic social media (Khalafat et al., 2021), Russian and Ukrainian Telegram (Nemkova et al., 2023), and English Twitter (Pilankar et al., 2022). Other tasks include threat detection for human rights defenders (Ran et al., 2023) and interpretable forced labor identification (Guzman et al., 2024). LLMs show potential for zero- and few-shot annotation but require human oversight (Nemkova et al., 2025b).
Conflict Prediction
NLP in conflict prediction mainly involves extracting and forecasting conflict dynamics using structured datasets like UCDP and ACLED (Hegre et al., 2019). Tasks include topic modeling for early signals (Mueller et al., 2024), modeling conflict stages with LLMs (Croicu and von der Maase, 2025; Nemkova et al., 2025a), automated report generation via retrieval-augmented generation (Nemkova et al., 2025c), and conflict content extraction and classification with domain-adapted models like ConfliBERT (Hu et al., 2022).
Physical Safety
NLP systems support understanding and intervention in threats to physical safety. They can be used to analyze social media, police reports, surveillance footage, health records, and news articles. Techniques like topic modeling More and Francis (2021), sentiment analysis Blevins et al. (2016), entity extraction Pavlick et al. (2016), and document classification Chang et al. (2018) help identify patterns of violence, perpetrators, and victims, informing interventions and policy decisions.
Challenges:
First, Data and Domain Complexity: references to violations and threats are often indirect or euphemistic, complicating detection. Labeled multilingual datasets—especially for low-resource or sensitive regions—are scarce ALSaif and Alotaibi (2019); MacPhaul et al. (2023); Parker (2020); Chang et al. (2018), and causes are often multifaceted and extend beyond textual signals. Annotation is labor-intensive with expert disagreement common Blandfort et al. (2019); Levy et al. (2022); Botelle et al. (2022); Chew et al. (2023), while evaluation suffers from the absence of standardized benchmarks and reliable ground truth, limiting model validation and comparison. Second, Model Limitations: NLP models struggle with code-switching, generalization across languages, regions, and out-of-distribution data, and often reflect training biases. The black-box nature of LLMs reduces transparency, hindering their deployment in high-stakes contexts, with real-world application still in early stages and lacking validated best practices. Lastly, Ethical and Evaluation Challenges: misclassifications can lead to legal or advocacy risks, highlighting the need for rigorous evaluation and ethical care.
Opportunities:
On a computational level, tools such as conflict prediction models and systems for detecting human rights violations can support evidence-based policymaking, inform the development and targeting of interventions, and enable rapid programmatic adjustments in response to evolving conditions. Developing multilingual and cross-regional models will enable more comprehensive monitoring. Further, NLP methods can contribute to operational efficiency for implementing organizations. For example, LLMs offer capabilities for rapid document review and synthesis, facilitating timely and contextually informed analysis. RAG systems can further enhance information access by efficiently extracting relevant content from document repositories. Additionally, LLMs can support scenario generation, situational analysis, and strategic planning. Lastly, a dedicated peacebuilding workshop could offer a crucial forum for interdisciplinary collaboration on NLP tools for humanitarian aid and crisis response.
6 [Uncaptioned image] Environment Protection
NLP offers scalable tools for climate mitigation and adaptation—supporting key SDGs (SDG6, SDG11, SDG12, SDG14, SDG15) and addressing critical global environmental risks (R2)—by extracting insights from unstructured text like scientific papers, policy reports, and assessments (on Climate Change, 2022). Techniques such as topic modeling (Sietsma et al., 2023), summarization (Ghinassi et al., 2024), and classification (Varini et al., 2020; Stammbach et al., 2023; Bingler et al., 2022; Schimanski et al., 2023) enable analysis of diverse datasets. NLP can also help to detect misinformation and greenwashing by verifying claims (Diggelmann et al., 2020; Hsu et al., 2024). While fine-tuning or pretraining LLMs on climate text is common (Leippold et al., 2022; Thulke et al., 2024), RAG-based chatbots (Vaghefi et al., 2023) and automated fact-checkers (Leippold et al., 2025) are emerging rapidly.
Challenges:
Extracting quantitative information, particularly from sustainability reports, remains challenging due to the complex, multi-modal and non-standardized nature of the data sources, leading to pipelines specifically designed for information extraction from tables (Mishra et al., 2024; Dimmelmeier et al., 2024). Furthermore, LLMs are prone to hallucination (Vaghefi et al., 2023) or are contaminated with false, conflicting or outdated climate information (Fore et al., 2024)Bulian et al. (2024) propose an evaluation framework based on presentational and epistemological qualities to evaluate quality and factual accuracy of LLMs.
Opportunities:
NLP is enabling cross-disciplinary collaboration, bridging domain experts with the fields of computer science, social science, and economics to analyze climate policies (Gandhi et al., 2024) or bringing the scientific community and non-governmental organization together to uncover narratives in public climate discourse (Gehring and Grigoletto, 2023; Rowlands et al., 2024). This is particularly important in expanding research to include cross-cultural and multilingual perspectives (Zhou et al., 2024; Bird et al., 2024).
7 [Uncaptioned image] Inclusion and Inequalities
Inclusive NLP that accounts for demographic, linguistic, and accessibility diversity is key to reducing systemic inequalities. These efforts support SDG5 and SDG10, and address inequality—a top global concern ranked among the most severe short- and long-term risks (GR4).
Gender Bias
Gender bias in NLP—especially in hiring De-Arteaga et al. (2019), healthcare, and moderation—can reinforce stereotypes and harm marginalized groups. Initial work focused on static embeddings using gender direction Bolukbasi et al. (2016) and WEAT Caliskan et al. (2017), later extending to contextual tasks like coreference Zhao et al. (2018) and occupation classification De-Arteaga et al. (2019). Recent studies examine LLM-agent interactions Borah and Mihalcea (2024) to track how biases spread. Mitigation includes post-processing (e.g., fine-tuning, DPO), interpretability, causal inference Attanasio et al. (2023); Cai et al. (2024), and counterfactual or intervention-based prompting Plyler and Chi (2025). Bias is commonly assessed via fairness gaps (e.g., F1, error disparities) or associations (e.g., WEAT, Bias-Score).
Cultural and Linguistic Representation
NLP systems often fail to capture the full diversity of human cultures and languages. Culture’s complexity - including evolving norms, values, and worldviews - varies widely across communities (Saha et al., 2025). Despite culture’s crucial role in communication, AI models frequently reflect incomplete cultural perspectives, overrepresenting dominant or cultural linguistic narratives typically from high-resource languages or communities (Hershcovich et al., 2022a; Mihalcea et al., 2025).
Underrepresented Communities
Users with diverse accessibility needs remain underserved, continuing to face significant communication barriers Khanuja et al. (2023a). NLP has been used in a wide range of assistive applications, including as augmentative communication Park et al. (2022), text simplification Espinosa-Zaragoza et al. (2023), text-to-speech Kumar et al. (2023), speech recognition Li et al. (2022), braille processing Tejesh et al. (2025), image captioning and subtitling Stefanini et al. (2021), question answering Gurari et al. (2018), assistive chatbots Grassini et al. (2024), reading aids Wang et al. (2024d), and sign language translation Rust et al. (2024).
Challenges:
Many issues revolve around binary gender assumptions and limited cross-linguistic/cultural generalizability Stanovsky et al. (2019); Adilazuarda (2024). Western-centric resources and lack of intersectional frameworks reinforce biases and marginalization  Kleinberg and Raghavan (2021); Sewunetie et al. (2024). Implicit biases and compounded discrimination, e.g., at the intersection of race, gender and disability, are underexplored Stanczak and Augenstein (2021); Guo and Caliskan (2021); Wald (2021). Data challenges persist due to scarce multimodal datasets, especially for Braille and sign language Hutchinson and Prabhakaran (2020); De Sisto et al. (2022). Limited interdisciplinary collaboration slows progress Kusters et al. (2020). Reliance on synthetic or prompt-generated data hampers real-world validity and hides systemic biases Venkit et al. (2025); Morales et al. (2024), while opaque data sourcing blocks accountability Bender and Friedman (2018).
Opportunities:
Inclusivity in NLP advances through participatory approaches centering marginalized voices, including co-designed fairness goals in gender bias mitigation Borah and Mihalcea (2024); Ma et al. (2023); Lauscher et al. (2022) and collaborative data collection with cultural experts and underrepresented groups Bird and Yibarbuk (2024); Newman-Griffis et al. (2024); UNICEF (2020); Hirmer et al. (2021). Solutions such as dynamic audit pipelines Park et al. (2023), lightweight model editing Park et al. (2023); Cai et al. (2024) and counterfactual data augmentation Zmigrod et al. (2019) help adapt models to sociocultural shifts. Fine-tuning LLMs can improve fairness and relevance Mai and Carson-Berndsen (2024); Bartl and Leavy (2024). Future work should consider identity compositionality Welch et al. (2020) and pluralistic alignment reflecting complex social affiliations Sorensen et al. (2024). Personalized, multimodal interaction design is vital for adaptive, accessible systems Paice et al. (2025); Wang et al. (2024d).
8 [Uncaptioned image]  Online Harms
The recent ease of access to digital devices (like smartphones and those based on IoT) has fueled the spread of digital violence globally (Bjelajac and Filipović, 2021). This is central to Technological Risks, one of the most critical risks identified in the 2025 Global Risks Report (GR5). There has been a large body of work on abusive/offensive/toxic/harmful speech classification Diaz-Garcia and Carvalho (2025), generation of counter speech Bonaldi et al. (2024); Saha et al. (2024); Wang et al. (2024a) and text detoxification Dementieva et al. (2025); Dale et al. (2021), including several languages Aluru et al. (2020).
Challenges:
Recent social media platform statements [^5] https://about.fb.com/news/2025/01/meta-more-speech-fewer-mistakes highlight the limitations of AI-based content moderation at scale. Challenges include the subjective nature of moderation, regional regulations, and the culturally diverse and implicit nature of content (Ocampo et al., 2023). LLMs struggle with volatile topics without frequent fine-tuning (Roy et al., 2023), and efforts are hindered by unclear label taxonomies, biases in pre-trained models, and limited collaboration between lawmakers, platforms, and researchers (Yimam et al., 2024).
Opportunities:
Opportunities for future work include building low-resource, robust, and generalizable LLM moderation frameworks that are deployable in real time, requiring stronger collaboration between researchers, moderators, and policymakers (Munzert et al., 2025; Bui et al., 2025). A clearer, more widely accepted taxonomy of digital violence is needed to reduce ambiguity in label definitions (Yimam et al., 2024). The current focus on text-based hate overlooks multimodal content like video and audio, calling for LLMs with multimodal moderation capabilities. To address culturally diverse and implicit content, future systems should integrate explainability, fact-checking, and real-time tools such as RAGs and search APIs. Finally, moderation efforts must expand beyond hate speech to cover the full spectrum of digital violence, while also embracing more inclusive approaches that consider author and receiver demographics, and broader cultural and linguistic diversity.
9 [Uncaptioned image] Misinformation
Ranked among the top five global risks in both short- and long-term scenarios (GR5), misinformation has a direct impact on the success of many of the SDGs. [^6] https://www.un.org/sites/un2.un.org/files/information-integrity-and-sdgs-en.pdf NLP approaches to misinformation include tasks such as fake news detection, rumor classification, stance detection, and fact checking (Oshikawa et al., 2020; Nakashole and Mitchell, 2014). Early methods used stylometric features, while modern systems use neural models, pre-trained LMs, and techniques like retrieval-augmented generation and prompt-based learning in multimodal and multilingual settings (Akhtar et al., 2023; Chen and Shu, 2024b).
Challenges:
There is a data scarcity, especially in low-resource languages, emerging domains, and shifting distributions (Guo et al., 2022). Multilingual and conflicting evidence complicates verification (Schlichtkrull, 2024; Zhang et al., 2024), while low-visibility claims targeting marginalized groups often go unchecked (Guo et al., 2022). In high-stakes domains like healthcare and politics, reliable detection is critical to prevent real-world harm (Abdul-Mageed et al., 2021; Zhao et al., 2023b), yet systems still lack robustness, fairness, and explainability. Lastly, LLMs pose misuse risks by generating convincing falsehoods (Buchanan et al., 2021; Gabriel et al., 2024).
Opportunities:
A key opportunity lies in developing human-centered evaluation methods tailored to real-world misinformation detection tasks (Das et al., 2023a). This would also benefit from interdisciplinary collaboration—for example, by integrating fundamental theories from social sciences and economics (Zhou and Zafarani, 2020), or by partnering with moderators, policymakers, and professional fact-checkers (Warren et al., 2025). Additionally, advancing the timeliness of ground truth through LLM-agents with access to web evidence and external knowledge bases can improve adaptability to fast-evolving misinformation. Developing domain-agnostic features that generalize across topics, languages, modalities, and time would help track shifting deceptive styles (Chen and Shu, 2024b) and identify check-worthy claims.
10 [Uncaptioned image] AI Harms
AI harms are among the top global threats (GR5). NLP is part of the problem—but also part of the solution. We start with the black-box problem and follow the harm taxonomy by Weidinger et al. (2022) to explore NLP mitigation strategies.
The Black Box Problem
LLMs often operate as black boxes, offering limited interpretability and transparency (Hassija et al., 2024). With little insight into their design or training, the public and AI community remain reliant on what creators choose to disclose. To address this, NLP research has explored textual explanations, such as summaries Atanasova et al. (2020); Kotonya and Toni (2020b), contrastive Schuster et al. (2021), and counterfactual forms Yang et al. (2020); Tolkachev et al. (2022). Visual methods like LIME (Ribeiro et al., 2016), ACE (Ghorbani et al., 2019), and heatmaps (Arras et al., 2017) highlight input relevance, while structural tools like SVCCA (Raghu et al., 2017), t-SNE, and TCAV (Kim et al., 2018) uncover concept-level insights.
Representation and Toxicity
NLP techniques are actively developed to address biases in training data and AI models (Gallegos et al., 2024), to reduce bias and toxicity. Further discussion is provided in §7 on inequality and §8 on digital toxicity.
Privacy, Safety and Malicous Uses
NLP approaches to harm reduction include learning to refuse or prevent memorization to protect personal data (Carlini et al., 2021; Liu et al., 2025b), adversarial training (Goyal et al., 2023), safe decoding (Xu et al., 2024b), authorship verification (Huang et al., 2024a), safety alignment (Bhardwaj et al., 2024), and red-teaming to expose vulnerabilities (Purpura et al., 2025).
Ungrounded Knowledge
As pointed in §9 LLM outputs can include hallucinations and spread misinformation (Pan et al., 2023; Huang et al., 2024b; Chen and Shu, 2024b). Mitigation strategies include grounding with external knowledge via RAG (Lewis et al., 2020; Peng et al., 2023; Asai et al., 2024), expressing uncertainty (Yang et al., 2023c; Feng et al., 2024; Xiong et al., 2024; Deng et al., 2024), and self-verification methods (Weng et al., 2023; Hong et al., 2024).
Socioeconomic and Environmental Harms
In NLP, researchers have called for greater climate awareness and transparency through reporting frameworks (Hershcovich et al., 2022b), while others focus on reducing energy use with model optimization techniques like pruning, quantization, and distillation (Schwartz et al., 2020; Jin et al., 2024; Zhu et al., 2024). Tackling socioeconomic harms (§4, §7) further requires closer collaboration with sociology and HCI (Blodgett et al., 2024; Card et al., 2024) to further enhance the limited work in this direction.
Challenges:
Key barriers to prevent AI harms include transitioning from prototype to deployment which entails coping with data distributional shifts between “lab” and “field” data, and managing bias or subjective labels. Many current benchmarks often overlook low-resource or politically sensitive domains (Kim et al., 2025; Cho et al., 2025). Additionally, the lack of standardized practices for measuring real-world impact complicates evaluation beyond standard metrics. The effectiveness of the proposed tasks is often questioned  (Chen and Shu, 2024a; Kotonya and Toni, 2024), and there are very few rigorous and holistic evaluation frameworks Atanasova et al. (2023); Liang et al. (2023). There are also several governance problems and establishing equitable partnerships with nonprofits, where power imbalances or misaligned goals can hinder collaboration.
Opportunities:
Field trials of NLP systems under real-world conditions are necessary, especially in low-resource or use-case-specific settings, while encouraging participatory design and critical user engagement (Pataranutaporn et al., 2025). On the modeling side, advances in retrieval-augmented generation (RAG) and knowledge augmentation methods can improve reliability by enhancing credibility (Xu et al., 2024a; Chen et al., 2024, 2025). Multidisciplinary approaches such as logical relationships between inputs (Ayoobi et al., 2025; Freedman et al., 2025) and mechanistic interpretations of model behavior (Hou et al., 2023; Yu and Ananiadou, 2024) can further enhance transparency. New research directions could be implemented by studying LLM overreliance, manipulation and unfair distribution of benefits from model access. Embracing process-aware NLP enables alignment with domain logic and ethical goals through explainability and fairness-by-design (Bernardi et al., 2024; Zhuang et al., 2025). Finally, mapping existing frameworks for AI4SG and NLP4PI (Fairness, Accountability, and Transparency in Machine Learning (2020), FAT/ML; Floridi et al., 2021), suggested actions NIST (2024) and principles for trustworthy AI (OECD, 2024), can support the development of adaptable, unified guidelines for responsible NLP deployment.
11 Summary and Call to Action
To achieve the full potential of NLP4SG, the community should move beyond task-centric innovation toward impact-driven systems that are safe, inclusive, and globally relevant. This work reviewed nine key application areas, addressing RQ1 on existing impactful solutions with concrete examples of how modern NLP technologies have already contributed. At the same time, in addressing RQ2 on key open challenges, we find that areas like misinformation, online harms, and education have seen sustained attention in NLP, while domains such as poverty alleviation, environmental protection, peacebuilding are only starting to gain traction in response to real-world crises. In response to RQ3 concerning promising directions for future work, we outline a set of universal challenges, opportunities, and actionable recommendations to guide future research in NLP4SG:
Recurring Challenges: Across these diverse domains, we identified five systemic challenges that continue to hinder progress: (1) Data scarcity and representational bias, particularly affecting low-resource languages and marginalized groups; (2) Misaligned evaluation metrics that fail to capture human-centered qualities like empathy or cultural sensitivity; (3) Safety, privacy, and ethical concerns, which are magnified in high-stakes settings such as mental health or misinformation detection; (4) Limited generalization across cultures, modalities, and real-world conditions; and (5) Infrastructural gaps and institutional fragmentation, especially in underrepresented scenarios.
Emerging Opportunities: Despite these barriers, we highlight several promising cross-cutting directions: (1) Multimodal and multilingual learning can help systems better reflect real-world diversity; (2) Human-AI collaboration enables more aligned, adaptive and interpretable NLP pipelines; (3) Participatory design and evaluation ensure that benchmarks and systems are co-developed with affected communities; (4) Retrieval-augmented and policy-aware methods offer new tools for building verifiable and context-sensitive applications; and (5) Explainability and AI literacy foster critical engagement and equitable access.
Call to Action: To advance NLP4SG, we call on the community to: (1) develop joint benchmarks featuring multilingual, culturally diverse, and socially grounded data; (2) collaborate closely with domain experts, such as educators, health practitioners, and civil society organizations, to co-design evaluation frameworks that reflect end-user needs; (3) pursue human-centered methodologies instead of one-size-fits-all solutions. Progress depends on pluralistic, context-aware roadmaps that align with both local realities and global development goals. Finally, while modern LLMs offer significant potential, it is crucial to ensure their affordability and accessibility so they serve the public good rather than exacerbate existing inequalities.
NLP has the tools to move beyond abstract benchmarking and toward socially responsive technologies designed with—and for—impacted communities. Realizing this vision requires not just technical innovation but also sustained interdisciplinary collaboration, inclusive practices, and a commitment to long-term global equity. We hope our findings can help earlier career researchers find their research niche and that more advanced researchers will have a fresh overview of the field to foster NLP4SG applications with a more interdisciplinary paradigm.
Limitations
This paper offers a high-level interdisciplinary perspective on aligning NLP research with societal needs, grounded in the UN Sustainable Development Goals (SDGs) and the World Economic Forum’s Global Risks Report. While our proposed framework maps NLP research directions to these agendas, the related works list discussed is not exhaustive. While we aimed to cover the most impactful topics based on the authors’ expertise, we acknowledge that ongoing advancements in NLP may enable new tasks and uncover deeper layers of impact beyond what could be envisioned at the time of writing.
Furthermore, while we highlight areas of overlap between the SDGs and global risks, the mappings remain somewhat subjective. Another key assumption we made is that positive impact is highly aligned with the UN Sustainable Development Goals - but this might not be true i.e. positive impact could still be derived from NLP tools without there being an explicit alignment towards SDGs. At the same time, we believe that our work will open more cross-disciplinary discussions and more ideas for NLP4SG applications.
Ethics Statement
This work is grounded in the belief that NLP research should be aligned with broader societal priorities and developed with care for its downstream impact. In proposing mappings between NLP directions, global goals, and risks, we are mindful of the potential for unintended consequences and overgeneralization. The open areas of work mentioned throughout the paper are just suggestive, and the practitioners, when working on them, should carefully evaluate the setting and downstream impacts.
We emphasize the importance of interdisciplinary collaboration and the inclusion of marginalized perspectives in shaping responsible NLP research. Where possible, we cite and build on existing initiatives in NLP for social good, and we support efforts to foreground equity, inclusivity, and transparency in both research framing and methodology. Finally, we acknowledge that AI assistants were used for proofreading this paper, a practice we consider ethically acceptable under appropriate usage.
Acknowledgments
This work was supported by the Novo Nordisk Foundation (grant NNF20SA0066568), which funded the research of Anders Søgaard and Antonia Karamolegkou. Antonia Karamolegkou was also supported by the Onassis Foundation - Scholarship ID: F ZP 017-2/2022-2023. Zhijing Jin was supported partly by the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039B; by the Machine Learning Cluster of Excellence, EXC number 2064/1 – Project number 390727645; and by Schmidt Sciences SAFE-AI Grant. Daryna Dementieva’s work was additionally supported by the Friedrich Schiedel TUM Fellowship.
Appendix A Paper Selection: Thematic Tables of Datasets, NLP Tasks, Evaluation Metrics, and References
A.1 Health and Well-being
Mental health is a key component of human health, encompassing our emotional, psychological, and social well-being SAMHSA (2023). Currently, mental health issues are a multifactorial global crisis complicated by individual risk factors and various socioeconomic and clinical factors, but NLP methods show promising potential to enhance mental healthcare Zhang et al. (2022). In this context, we define NLP tasks as activities performed by NLP techniques in roles similar to counselors and clients. In the role of counselors, NLP tools engage in several core tasks: (1) Detection and classification of mental health conditions, such as depression Giuntini et al. (2020); Mahdy et al. (2020); Khan et al. (2018) and addiction Yang et al. (2023a); Kwon et al. (2023); Ni et al. (2021), using data sources like clinical notes Panaite et al. (2022); Calvo et al. (2017) and social media posts Skaik and Inkpen (2020); Chancellor and De Choudhury (2020); Ríssola et al. (2021); (2) Responding to users by interpreting their emotional states Warikoo et al. (2022); Sabour et al. (2024); Grandi et al. (2024), generating therapeutic and empathetic responses Shen et al. (2020); Grandi et al. (2024); Sharma et al. (2023); Nazarova (2023); Zhou et al. (2025), and providing actionable feedback for support quality Min et al. (2023); Chaszczewicz et al. (2024); Althoff et al. (2016); (3) Tracking emotion and mood via time-series data analysis Čosić et al. (2024) and detecting mental health crises over time Gong et al. (2019); Yuan et al. (2023). Conversely, when NLP tools serve as clients, they typically simulate client personas from diverse backgrounds to train counselors Louie et al. (2024); Liu et al. (2025a); Hsu et al. (2025); Kampman et al. (2025). The literature for this section was selected based on the existing survey papers such as Zhang et al. (2022); Malgaroli et al. (2023) as well as keyword search (e.g., "AI for mental health", "LLM-based mental health applications", "mental health chatbots", "AI therapy").
Physical well-being refers to maintaining one’s bodily health through behaviors such as regular physical activity, adequate sleep, balanced nutrition, good hygiene, and avoidance of harmful substances Lotan et al. (2005). To stimulate discussion on the role of NLP in physical well-being, the works included in the paper were selected based on a targeted keyword search including “physical activity”, “sleep”, “nutrition”, “hygiene”, “substance use”, and “clinical report analysis”, which also plays a central role in physical health. NLP techniques have been applied to various aspects of physical well-being, leveraging unstructured text data to monitor behaviors and inform interventions. For instance, physical activity and sedentary behavior can be tracked through analysis of social media using NLP-driven health surveillance systems that mine Twitter posts to estimate physical activity levels, sedentary behavior, and sleep patterns in populations Sakib et al. (2021); Shakeri Hossein Abad et al. (2022). For nutrition, NLP has been employed to assess dietary habits by using models that can automatically classify foods and meals from descriptions, even providing personalized diet advice van Erp et al. (2021); Hu et al. (2023). Regarding hygiene, during the COVID-19 pandemic NLP was used to estimate public perceptions of mask-wearing and other hygiene practices by mining social media posts Al-Garadi et al. (2022b). Likewise, in harmful habit avoidance, NLP methods help identify substance use patterns and risky behaviors from online text Hu et al. (2021); Lin et al. (2023). The analysis of clinical reports also plays a central role in physical health. These documents capture critical patient information that is often not represented in structured data fields. Consequently, clinical report analysis has emerged as a core subtask within NLP for physical health, enabling the extraction, classification, and summarization of medically relevant information directly from unstructured clinical narratives Landolsi et al. (2023). In this research area there is a growing role of LLMs in extracting information from clinical reports. The work by Mannhardt et al. (2024) shows how GPT-4 can support patients by simplifying clinical notes, improving comprehension and confidence, though with some factual inaccuracies. In Guluzade et al. (2025), the authors introduce a large annotated dataset (ELMTEX) and find that fine-tuned small LLMs outperform larger ones in extracting structured information efficiently. These findings are confirmed in pathology reports, where fine-tuned models achieve higher accuracy and fewer hallucinations than prompt-based methods Park et al. (2024). LLMs can also be used in clinical reports to generate and summarize documentation such as patient notes, discharge summaries, and case reports, offering improvements in efficiency, organization, and standardization of medical writing Park et al. (2024); Ali et al. (2023); Patel and Lam (2023); Cascella et al. (2023). They can help identify grammar errors and inconsistencies in extracted data (e.g., lab values), thereby potentially reducing documentation errors Ali et al. (2023). These applications may alleviate administrative burdens on healthcare professionals, allowing more time for direct patient care Lee et al. (2023). Nonetheless, the performance of LLMs is limited by variability in accuracy depending on case complexity, the risk of generating incorrect or fabricated content (hallucinations), and susceptibility to user framing, emphasizing the need for careful prompt design and human oversight Puthenpura et al. (2023).
Table 1: Overview of datasets, NLP tasks, evaluation metrics, and references from healthcare-related studies.
Datasets NLP Task(s) Evaluation Metrics Reference
Expert-annotated opioid-related posts on Reddit Detecting addiction Accuracy, macro-F1 Yang et al. (2023a)
Motivational Interviewing (MI) dataset Generating therapeutic dialogues ROUGE, embedding-based metrics (greedy matching, embedding average, and vector extrema), ratio of distinct n-grams, human annotator evaluation Shen et al. (2020)
Empathetic Dialogues Dataset, Reddit Mental Health Dataset, DailyDialog Dataset Generating empathetic dialogues for mental health support BERTSCORE, accuracy, precision, recall, F-1 Grandi et al. (2024)
MI-TAGS Publicly available mental health dataset Accuracy, macro F-1, ROC AUC Cohen et al. (2024)
Reddit Self-reported Depression Diagno- sis dataset Interpretable NLP models in mental health Precision, recall, F-1 Song et al. (2018)
Depression dataset, Non-depression dataset, Depression-candidate dataset Interpretable NLP models in mental health Precision, recall, F1, accuracy Zogan et al. (2021)
CAMS Causal Analysis of Mental health issue Accuracy Garg et al. (2022)
Dataset of principles Simulating patient personas Consistency with Context, speech style, Principle Adherence Louie et al. (2024)
Publicly available depression-related conversations (RED, HOPE, ESC, AnnoMI-Full), expert-annotated preferences Simulating patient personas Expert evaluation on contrast with AI-like responses, linguistic authenticity, cognitive pattern authenticity, subtle emotional expression, profile adherence and personalization. Automatic evaluation on symptom severity, cognitive distortion, and overall depression Severity Liu et al. (2025a)
MIDAS Publicly available dataset in mental health counseling Expert evaluation, reflection to question ratio, accuracy, F-1 Gunal et al. (2025)
Reddit Mental Health Dataset Publicly available mental health dataset on social media Recall, precision, F-1 Wu et al. (2022)
Longitudinal Patient Health Questionnaire Tracking user mood or mental health crises Spearman’s rank-order correlation, mean squared error Gong et al. (2019)
FeedbackESConv Providing feedback to counselors Automatically-computed quality scores, domain experts Chaszczewicz et al. (2024)
PAIR, AnnoMI Providing feedback to counselors Edit effect (reflection score), content preservation, perplexity, coherence, specificity Min et al. (2023)
Anonymized counseling conversations from a NGO Providing feedback to counselors Adaptability, dealing with ambiguity, creativity, making progress, change in perspective Althoff et al. (2016)
Annotated clinical notes Predicting and understanding mental health outcomes Accuracy Panaite et al. (2022)
Thought Records Dataset, Mental Health America Responding to users’ negative thoughts Automatic (BLEU, ROUGE-1, ROUGE-L, BertScore); Human (Relatability, Helpfulness)
Insomnia data set consisting from Twitter Text classification, Correlation analysis betwen language use and insomnia, Topic modeling True-positive rate, False-positive rate, AUC Sakib et al. (2021)
Twitter corpus (LPHEADA) labeled for relevance to physical activity, sedentary behavior, and sleep Text classification, Semantic consistency evaluation, location inference Precision, Recall, F1 Score, AUC-ROC, Average precision Shakeri Hossein Abad et al. (2022)
Various online recipe databases, both structured and unstructured: recipe websites, historical recipe archives, nutritional databases, sustainability data NER, Information extraction, semantic linking, recommender system Qualitative analysis van Erp et al. (2021)
Food Label Information and Price (FLIP) Database Text classification, Regression, t-SNE visualization Accuracy, Precision, Recall, F1-score, MSE Hu et al. (2023)
Electronic health records, social media platforms (Twitter, Reddit, Facebook, YouTube), scientific literature, news and web sources Information Extraction, Health Behavior Analysis, Early outbreak detection, Misinformation detection, Question Answering Accuracy, Precision, Recall, F1-score (for classification), AUC-ROC, MSE (for regression) Al-Garadi et al. (2022b)
Posts manually labeled with six annotation categories from Reddit NER, WSD, Sequence labeling, Social media text analysis Precision, Recall, F1-score Hu et al. (2021)
Facebook posts from antitobacco campaigns Sentiment analysis, Topic modeling, Text classification Odds ratios, Agreement rate Lin et al. (2023)
Hypothetical clinical scenarios related to skin cancer Text generation, Readability assessment Readability score, Likert ratings Ali et al. (2023)
Simulation of real-world clinical and research use cases Clinical note generation, Detection of potential misuse, Language style adaptation Qualitative analysis Cascella et al. (2023)
Real-world pathology reports Information extraction, Hallucination detection Accuracy Park et al. (2024)
12 clinical notes: 4 synthetic and 8 real Text simplification, Definition extraction, FAQ generation, Information extraction, Prompt engineering Quantitative evaluation of a survey, Readability score, Qualitative interviews Mannhardt et al. (2024)
ELMTEX Dataset (clinical summaries) Information extraction, Entity normalization ROGUE, BERTScore, Precision, Recall, F1 Guluzade et al. (2025)
A.2 Education
NLP for Education Tools & Systems.
The integration of NLP systems into educational settings has garnered significant attention, particularly with the widespread adoption of LLMs by students. These NLP-based educational applications, such as intelligent tutoring systems, offer the potential to deliver personalized, high-quality education to underserved regions and populations. Specifically, these systems aim to enhance learning experiences by providing timely and personalized support to both teachers and students, including the following tasks: personalized and/or curriculum-aligned question generation Kargupta et al. (2024); Lucy et al. (2024), scaffolded dialogue tutoringKazemitabaar et al. (2024), adaptive knowledge tracing Kargupta et al. (2024), automated feedback Jurenka et al. (2024), teacher coaching Wang and Demszky (2023), and student simulation for testing classroom policies/activities Zhang et al. (2025).
Methodologies.
Intelligent tutoring systems primarily focused on (1) modeling teacher-student and student-student interactions using transcripts, (2) devising knowledge state spaces for specific domains/problems to trace student knowledge throughout an interaction, and (3) generating single-turn responses to students. With the emergence of LLMs, recent methodologies have expanded upon these tasks to include:
  1. 1.
    Multi-Turn Socratic Dialogue and Planning. Recent methodologies leverage large language models (LLMs) to engage students in multi-turn Socratic dialogues, promoting critical thinking and problem-solving without directly providing answers. For instance, TreeInstruct employs a state space-based planning algorithm to dynamically construct question trees based on student responses, effectively guiding learners through multi-turn code debugging tasks Kargupta et al. (2024). Similarly, the Socratic Questioning of Novice Debuggers dataset benchmarks LLMs’ abilities to employ Socratic methods in assisting novice programmers through single-turn interactions Al-Hossami et al. (2023).
  2. 2.
    Expert Decision Modeling. To emulate expert tutoring behaviors, some works model the decision-making processes of experienced educators. Bridging the Novice-Expert Gap utilizes cognitive task analysis to capture experts’ identification of student errors, remediation strategies, and instructional intentions, informing LLM responses in math tutoring scenarios Wang et al. (2024b).
  3. 3.
    Curriculum-Aligned Evaluation. Evaluating LLMs’ mathematical reasoning has shifted towards alignment with educational curricula. MathFish assesses whether models can identify and apply specific math skills and concepts as outlined in standardized curricula, using publisher-labeled data from open educational resources Lucy et al. (2024).
  4. 4.
    Open-Ended Pedagogical Benchmarking. To assess LLMs’ instructional capabilities beyond problem-solving, MathTutorBench introduces a benchmark evaluating open-ended pedagogical skills. It measures models’ abilities across various educational tasks, emphasizing the quality of instructional interactions Macina et al. (2025).
  5. 5.
    Simulated Student Interactions. Datasets like MathDial are created by pairing human teachers with LLMs simulating student behavior, generating rich pedagogical dialogues. This approach aids in training and evaluating models on realistic tutoring scenarios Macina et al. (2023).
  6. 6.
    Classroom Discourse Analysis. Large-scale datasets such as the NCTE Transcripts provide insights into teacher-student interactions. These transcripts, annotated for dialogic discourse moves, help in analyzing effective instructional practices and inform the development of NLP tools for education Demszky and Hill (2023).
  7. 7.
    Educational Conversation Toolkits. Open-source frameworks like Edu-ConvoKit facilitate the analysis of educational conversations by offering tools for preprocessing, annotation, and analysis tailored to educational research needs Wang and Demszky (2024).
We have included an overview of the various dataset resources and their corresponding tasks and evaluation metrics in Table LABEL:tab:edu_datasets. These papers have been collected based on recent prominence and relevance to different challenges and opportunities present within the NLP Education space.
Table 2: Overview of datasets, NLP tasks, evaluation metrics, and references from NLP for Education studies.
Datasets NLP Task(s) Evaluation Metrics Reference
MathDial Text-to-Text Generation (Tutoring Response Generation) sBLEU, BERTScore, KF1, Uptake, Success@k, Telling@k, Human Evaluation (Coherence, Correctness, Equitable tutoring) Macina et al. (2023)
MULTI-DEBUG Text-to-Text Generation, Text Classification (Socratic Question Generation, Multi-Turn Planning) Relevance, Indirectness, Logical Flow, Overall Success Rate, Average # of Turns Kargupta et al. (2024)
Bridge Text Classification, Text-to-Text Generation (Remediation of Math Mistakes, Decision-Making Modeling) Human Evaluation (usefulness, care, human-soundingness, preference), Log Odds Ratio Wang et al. (2024b)
MathFish Text Classification, Topic Modelling, Text-to-Text Generation (Math Reasoning Evaluation, Curriculum Alignment) Weak Accuracy, Exact Accuracy Lucy et al. (2024)
MathTutorBench Text-to-Text Generation (Evaluation of Pedagogical Capabilities in LLM Tutors) Accuracy, BLEU, F1, Win Rate (Pedagogical skill metrics) Macina et al. (2025)
National Center for Teacher Effectiveness (NCTE) Text Classification (Evaluation of Classifying Educational Discourse Features) Accuracy, Precision, Recall, F1 Demszky and Hill (2023)
A.2.1 AI Literacy
To identify relevant literature on AI literacy, we used Semantic Scholar and Google Scholar search terms such as “ai literacy” and (“ai literacy” + “social impact” + “nlp”). We found most of the papers selected for this topic to be on classroom-based studies and measurements and metrics for AI literacy, along with some interdisciplinary papers connecting AI literacy to other disciplines, for example, psychology.
A.3 Peace Building
We have organized highlighted papers found in our review of papers at the intersection of peace building and NLP in Table 3.
To identify the most relevant literature on human rights violation detection using NLP and on conflict prediction, we employed three search strategies: querying the ACL Anthology, conducting searches on Google Scholar, and utilizing the Consensus research discovery platform. The keywords used included: “human rights,” “human rights violations detection,” “armed conflict prediction,” and “conflict forecasting.”
The selected works on physical safety were identified through a keyword search in ACL and Google Scholar for papers on “physical safety”, “domestic violence”, “gun violence” and “firearm injury”. The Google Scholar search also included the keyword “nlp”. We did not include papers that were more focused on the mental health implications of physical safety (e.g. suicide via firearms), or on larger organizational peace-building efforts (e.g. terrorism and police brutality).
We find most of these tasks in this topic to be focused on document classification of rare events. As a result, primary evaluation metrics are precision, recall, and F1. There are a few papers that apply unsupervised tasks like topic modeling and generation that use coherence and similarity, respectively, as their primary metrics.
Datasets NLP Task(s) Evaluation Metrics Reference
Crisis Text Line Database Document Classification (detect firearm injury or violence) Precision, Recall, Accuracy Chew et al. (2023)
National Violent Death Report System Topic Modeling (characterize trends in violent deaths) Coherence, Topic diversity, Coverage Arseniev-Koehler et al. (2022)
National Electronic Injury Surveillance System Series Document Classification (classify location of nonfatal gunshot injuries) Accuracy, Precision, Recall, AUC Parker (2020)
SafeText (Reddit) Generation (generating advice) Similarity, Confidence, Perplexity, Accuracy Levy et al. (2022)
Surveillance Cameras Speech to Text, Document Classification (Detect violence) Precision, Recall, Accuracy, Loss Kumari et al. (2023)
Twitter Document Classification (detect violence related tweets) Precision, Recall, Accuracy ALSaif and Alotaibi (2019)
Police Reports Document Classification (Classify abuse type and victim injuries) Precision, Recall, F1 Karystianis et al. (2019)
Twitter Document Classification (detect intimate partner violence) Accuracy, F1 Al-Garadi et al. (2022a)
Electronic Health Records Document Classification (classify firearm injury intent) Precision, Recall, F1 MacPhaul et al. (2023)
Gun Violence Database (deprecated compilation of news articles) Entity Extraction (identify event details like participants, roles, location, time) Precision, Recall Pavlick et al. (2016)
Twitter POS tagging, Machine translation, Sentiment Analysis, Document Classification (detect aggression and loss) Precision, Recall, F1 Blevins et al. (2016)
Twitter Document Classification (detect aggression, loss, and substance use) Precision, Recall, F1, Average Precision Blandfort et al. (2019)
Twitter Document Classification (detect aggression and loss) Precision, Recall, F1 Chang et al. (2018)
Electronic Health Records Document Classification (If a patient will become violent) F1, Confusion Matrices Borger et al. (2022)
Electronic Health Record Document Classification (type of violence and patient status) Precision, Recall, F1 Botelle et al. (2022)
Table 3: Overview of Peace Building and Physical Safety related NLP studies
A.4 Poverty
We began our review of papers focused on the study of poverty by through keyword searches such as “poverty detection" on Google Scholar. However, the majority of these studies addressed this issue through the use of satellite imagery Tingzon et al. (2019); Ayush et al. (2020), or audience estimates from advertising platforms (Fatehkia et al., 2020). From these sets of papers, we included a review of the use of AI for poverty prediction Usmanova et al. (2022). We then modified our keyword searches on Google Scholar to “nlp income", “nlp poverty", and “text poverty classification" to ensure a focus on studies from the NLP domain. For papers related to our task, we would investigate studies that have cited them as well. We have highlighted a number of the most prominent papers found in our review in Table 4.
The current NLP literature on the topic of poverty and class is rather limited Cercas Curry et al. (2024). A literature review published in 2024, only found 20 NLP papers that investigate socio-economic status in any capacity Cercas Curry et al. (2024). Another review of artificial intelligence systems for the detection of poverty Usmanova et al. (2022), conducted in 2022, identified 22 papers, only one of which used NLP models Muñetón-Santa et al. (2022). Additionally, most of the datasets are not publicly available, hindering reproduction and progress in this domain.
Datasets NLP Task(s) Evaluation Metrics Reference
Twitter Yearly Income Prediction Pearson correlation, Mean Average Error Preoţiuc-Pietro et al. (2015b)
Twitter Income Prediction, Temporal Orientation Classification Accuracy, Precision, Recall, F1, MAE Hasanuzzaman et al. (2017)
Interview Transcripts Poor or Extremely Poor Classification Accuracy, Specificity, Sensitivity, F1 Muñetón-Santa et al. (2022)
News ESI Prediction, Unemployment Prediction RMSE Lampos et al. (2014)
Table 4: Overview of Poverty related NLP studies
A.5 Online Harms
Initial research primarily focused on developing classification frameworks across different spectrum of online harms based on existing AI models. Starting from traditional ML tools (Saha et al., 2018), further works exploited LSTMS/RNNs (Kumar, 2022), and with the advent of Transformers based models (like BERT Saleh et al. (2021)), research rapidly unfolded with the increased development of frameworks brewed on top of attention mechanism. Recent advancements of LLMs has lead to works using their generation capability Guo et al. (2024); Tassava et al. (2024); Pendzel et al. (2023); where some works also incorporate infamous LLM strategies like zero-shot (Roy et al., 2023) and few-shot Zahid et al. (2025) promptings, and fine-tuning Nasir et al. (2025). For explainability, (Mathew et al., 2022) presented one of the first works that proposed one-hot vector representation to improve attention based models and recently, reasoning based explanation and interpretation frameworks (Yang et al., 2023b; Nirmal et al., 2024) to provide more contextual information to LLMs have garnered attention.
Apart from traditional classification based mitigation, a rapid shift towards proactive content moderation leveraging the generative capabilities of LLMs has been proposed. Numerous works, especially focused on two strategies– counter speech generation Bonaldi et al. (2024); Saha et al. (2024); Wang et al. (2024a) and text detoxification Dementieva et al. (2025); Dale et al. (2021) have been extensively explored. LLMs have proven to be sufficiently good at these tasks, but need further improvements for multilingual performance Dementieva et al. (2024). Some recent works have further proposed the effectiveness of strong few-shot capabilities of LLMs for annotation of such complex datasets which can potentially reduce crowd-sourcing efforts (Bhat and Varma, 2023; Kim et al., 2024). Further, studies on curating multimodal datasets Kiela et al. (2021) and understanding the strengths and limitations of multimodal LLMs have also garnered attention Rizwan et al. (2025). We have highlighted significantly important datasets and studies on online harms in Table 5. These papers are shortlisted on the basis of the impact they drive thus aiding in the detection and mitigation of online harms through necessary content moderation strategies.
Datasets NLP Task(s) Evaluation Metrics Reference
Twitter Multilingual and Multi-Aspect Hate speech detection spanning different target communities Micro-F1 and Macro-F1 score Ousidhoum et al. (2019)
Twitter + Gab (HateXplain) Hate, Offensive and Normal speech detection with rationales Classification - Accuracy, Macro-F1 score, AUROC score; Rationales - IOU-F1 score, token-F1 score, AUPRC score (Plausibility) and comprehensiveness, sufficiency (Faithfulness) Mathew et al. (2022)
LLM generated explanations Explainable hate speech detection with step-by-step reasoning generated by LLMs Accuracy and F1 score Yang et al. (2023b)
Hate-COT Offensive speech label explanation generated by GPT-3.5 Turbo F1 score, Persuasiveness and Soundness Nghiem and Daumé Iii (2024)
Latent Hatered Detection of implicit hate speech Precision, Recall, F1 score and Accuracy ElSherief et al. (2021)
Jigsaw toxicity datasets Detection of different types of toxicity across multiple labels with corresponding severity score Overall AUROC and Bias AUROC cjadams et al. (2017, 2019); Kivlichan et al. (2020)
Measuring hate speech (Comments from YouTube, Reddit and Twitter) Rasch Measurement Theory (RMT) based continuous scoring of hate speech across multiple labels and targets Hate speech score, difficulty of survey item and response, severity of rater Sachdeva et al. (2022)
CONAN and its variants Generation of counter speech against hate speech through different NLP generation strategies Semantic Similarity, Novelty, Diversity, Toxicity, Politeness, Intent Accuracy and Hate Mitigation Chung et al. (2019); Fanton et al. (2021); Bonaldi et al. (2022); Gupta et al. (2023)
Toxic instances from Jigsaw, Reddit and Twitter Toxicity classifier and generation of detoxified speech for toxic instances Accuracy, Fluency, Similarity and Joint score Logacheva et al. (2022)
Multilingual text detoxification dataset from multiple sources Multilingual text detoxification with explanation Style Transfer Accuracy (STA), Fluency (ChrF1), Content Similarity and Joint score Dementieva et al. (2025)
Facebook + Twitter memes Hate, harm and misogyny detection in memes using different multimodal applications of NLP Accuracy, Macro-F1 score and AUROC score Kiela et al. (2021); Pramanick et al. (2021); Fersini et al. (2022)
BitChute videos Hateful videos classification at the intersection of NLP, vision and audio Accuracy, Macro-F1 score, Precision and Recall Das et al. (2023b)
Table 5: Representative datasets, NLP task(s) and evaluation metrics from research on online harm.
A.6 Misinformation
Methodology. Automated fact-checking process comprises the verification of claims - verifiable factual statements (Panchendrarajan and Zubiaga, 2024). Four major components of the fact-checking pipeline (Das et al., 2023a) are widely studied (Vlachos and Riedel, 2014; Thorne and Vlachos, 2018; Barrón-Cedeño et al., 2020; Guo et al., 2022) and include: (1) claim detection, checkworthiness, and prioritisation (based on their urgency or potential harm / impact) to identify claims from news and social media that should be processed given the limited human and automated fact-checking resources (Konstantinovskiy et al., 2021; Nakov et al., 2021; Abumansour and Zubiaga, 2023), often treated as a classification task; (2) evidence retrieval to collect trustworthy evidence for a claim (Thorne et al., 2018a; Augenstein et al., 2019); (3) veracity prediction based on this evidence; and, finally, (4) explanation of the outcome label for humans (Shu et al., 2019; Kotonya and Toni, 2020a; Atanasova et al., 2020; Lu and Li, 2020): summarizing the evidence, generating explanations and evaluating them. The existing datasets align with the fact-checking stages. They are included in Table 6. Other close tasks from the automated misinformation detection field, such as stance detection, rumor detection and fake news detection based on linguistic features are also included in this Table. In addition to this, there also more domain-specific datasets for misinformation detection - for example, multiple datasets were collected on COVID-19 topic (Abdul-Mageed et al., 2021; Song et al., 2021; Mohr et al., 2022; Heinrich et al., 2024), including a multilingual dataset for a shared task to predict fact-checking options for claims, including their verifiability and potential harm (Shaar et al., 2021).
Misinformation can be spread in various languages, but most datasets are still in English. Multilingual verification can use translation systems, but datasets in specific languages and multilingual datasets are still needed to train and evaluate monolingual and multilingual models - for example, comprehensive multilingual multitopic claim detection datasets (Panchendrarajan and Zubiaga, 2024) (despite the recent efforts e.g. in (Nakov et al., 2021; Kazemi et al., 2022a; Pikuliak et al., 2023)).
For our review, we firstly focused on surveys on automated misinformation detection (Guo et al., 2022; Das et al., 2023a; Panchendrarajan and Zubiaga, 2024; Khiabani and Zubiaga, 2024; Huang et al., 2025; Chen and Shu, 2024b; Oshikawa et al., 2020; Zhou and Zafarani, 2020), then studied papers and approaches mentioned.
Datasets NLP Task(s) Evaluation Metrics Reference
Emergent stance classification accuracy, per-class precision and recall (Ferreira and Vlachos, 2016)
Multi-Target Stance Dataset stance classification macro-averaged F1 score (Sobhani et al., 2017)
PHEME rumor detection and verification; stance classification macro F1 score, accuracy (Kochkina et al., 2018)
RumourEval 2019 stance towards a rumor: classification; veracity prediction: classification macro F1 score; macro F1 score, RMSE (Gorrell et al., 2019)
VAST stance classification macro-averaged F1 score (Allaway and McKeown, 2020)
Will-They-Won’t-They stance classification for rumor verification macro F1 score, unweighted avg F1, weighted avg F1 (Conforti et al., 2020b)
STANDER stance classification; evidence retrieval macro-averaged precision, recall and F1 score; precision@5 and recall@5 (Conforti et al., 2020a)
COVID-19-Stance classification accuracy, macro average precision, recall, F1 score (Glandt et al., 2021)
P-Stance stance classification F avg, macro-average of F1 score (Li et al., 2021)
ISD stance detection classification micro average F1 score (Huang et al., 2023)
C-STANCE stance classification F1 scores for 3 classes and and F1 macro (Zhao et al., 2023a)
MT-CSD stance classification F avg (Niu et al., 2024)
TSD-CT stance classification F1 scores for each class and macro F1 score (Zhu et al., 2025)
LIAR fake news: classification accuracy (Wang, 2017)
FakeNewsAMT and Celebrity fake news: classification accuracy, precision, recall, and F1 score (Pérez-Rosas et al., 2018)
Stance-annotated Reddit dataset rumor stance and veracity prediction: classification accuracy, F1 score (Lillie et al., 2019)
Twitter-based dataset classification accuracy, average precision, ROC, F1 micro, F1 macro scores (Volkova et al., 2017)
Claim detection dataset claim detection: classification precision, recall and F1 score (Konstantinovskiy et al., 2021)
MultiFC Claim verification: classification; evidence ranking micro F1, macro F1 (Augenstein et al., 2019)
Snopes-based dataset stance classification; evidence extraction: ranking; claim validation: classification precision, recall and F1 macro; precision @5 and recall @5; macro presicion, recall and F1 (Hanselowski et al., 2019)
CLIMATE-FEVER claim verification: retrieval, ranking and classification accuracy (Diggelmann et al., 2020)
SciFact claim verification: retrieval and classification precision, recall, F1 score (Wadden et al., 2020)
PUBHEALTH veracity prediction: classification; explanation generation precision, recall, F1 macro, accuracy; ROUGE and coherence (Kotonya and Toni, 2020b)
COVID-Fact evidence retrieval, claim verification: classification COVID-FEVER Score (similar to FEVER score) (Saakyan et al., 2021)
X-Fact claim verification: classification F1 score (Gupta and Srikumar, 2021)
FakeNewsNet claim verification: classification precision, recall, accuracy, F1 score (Shu et al., 2018)
FEVER claim verification: retrieval, ranking and classification accuracy, F1 score; FEVER score (includes evidence retrieval and claim labels) (Thorne et al., 2018a, b)
FEVEROUS claim verification: retrieval, ranking and classification FEVEROUS score (includes evidence retrieval and claim labels) (Aly et al., 2021)
Multilingual claim matching dataset claim matching: retrieval and classification MAP@k, MRR and F1 score, accuracy (Kazemi et al., 2022b)
CLEF-2022 CheckThat! Task 2 claim matching: ranking MAP, reciprocal rank, Precision@k, MAP@5 (Nakov et al., 2022)
MultiClaim claim matching: ranking S@10 (Pikuliak et al., 2023)
NLP4IF-2021 claim detection: classification precision, recall, F1 score (Shaar et al., 2021)
CLEF-2022 CheckThat! Task 1 verifiable claim detection: classification F1 score, accuracy, weighted F1 (Nakov et al., 2021)
CLEF-2024 CheckThat! Task 5 evidence retrieval; rumor classification MAP and Recall@5; F1 macro and strict F1 macro (Haouari et al., 2024)
Table 6: Datasets on misinformation detection.
A.7 Inequalities and Bias
Datasets NLP Task(s) Key Evaluation Metrics Reference
GloVe, Word2Vec embeddings on Google corpus Intrinsic bias probing WEAT, gender–direction cosine Bolukbasi et al. (2016); Caliskan et al. (2017)
WinoBias Coreference resolution F1 / precision gap (Female vs Male) Zhao et al. (2018)
WinoMT Machine translation (pronoun gender) Gender accuracy, error rate Stanovsky et al. (2019)
BiosBias (LinkedIn biographies) Occupation classification F1 diff., error disparity De-Arteaga et al. (2019)
LLM-Agent Interaction Logs Multi-agent bias propagation Bias-Score, fairness gap Borah and Mihalcea (2024)
BiosBias+GPT-J Causal weight-editing for debiasing Stereotype score, accuracy Cai et al. (2024)
TED-Talk / IWSLTEn–It MT Pronoun-drop attribution maps Pronoun-drop rate, TER gap Attanasio et al. (2023)
Synthetic counterfactual pairs (Iter CDA) Bias-robust text classification Bias amplification ratio, F1 Plyler and Chi (2025)
GeoWAC, Reddit, UN General Debates Region-aware bias evaluation metric Region-aware WEAT effect size, mismatchBorah et al. (2025)
CrowS-Pairs (intersectional ext.) Intrinsic bias evaluation Bias Score (Black vs White) Guo and Caliskan (2021)
Table 7: Representative datasets and evaluation practices across gender-bias NLP tasks. These were some influencial datasets and papers in gender bias in NLP, covering the topics of bias detection and bias mitigation methods across static embeddings, dynamic embeddings, transformer-based LMs, and LLMs
Datasets NLP Task(s) Key Evaluation Metrics Reference
GeoDE, GD-VCR, CVQA Multimodal Captioning; Multi-Agent Collaboration Alignment score; Completeness score; Cultural Info metric Bai et al. (2025)
XNLI, PAWS-X Cross-lingual NLI; Paraphrase Detection Per-culture accuracy gaps Hershcovich et al. (2022a)
World Values Survey (WVS-7) Survey-Response Prediction; Alignment Measurement Similarity scores; Alignment gap per group AlKhamissi et al. (2024)
CultureBank TikTok, CultureBank Reddit Cultural QA; Zero-Shot QA; Fine-Tuning QA accuracy improvements; Agreement levels Shi et al. (2024)
NormAd-Eti Norm Classification (acceptable vs not) Model accuracy vs human on explicit/abstract norms Rao et al. (2025)
GD-VCR Culturally-Aware Captioning Human eval (cultural descriptiveness ratings) Yun and Kim (2024)
C4 web crawl Commonsense Extraction; Classification; Clustering Crowdsourced plausibility (PLA, COM, DIS); QA priming gains Nguyen et al. (2023)
Dollar Street Zero-Shot Image–Text Alignment Median CLIP score by income quartile; Spearman $\rho$ Nwatu et al. (2023)
Universal Dependencies treebanks; SIGMORPHON; WMT news translation; XNLI cross-lingual NLI; TyDi QA/ SQuAD Parsing; Inflection; MT; TTS; NLI; QA Scaled performance utility; Global utility metrics Blasi et al. (2021)
ImageNet Image Classification by Country Accuracy gaps (US/EU vs developing regions) Shankar et al. (2017)
WikiAnn; Universal Dependencies treebanks; XNLI; TyDI QA/ChAII NER; POS; NLI; QA Utility × Demand; Gini coefficient; Throughput and memory Khanuja et al. (2023b)
Google Street View images; American Community Survey data; Voting precinct results Vehicle Detection; Attribute Classification; Demographic Regression Correlation vs ACS; Voting prediction accuracy Gebru et al. (2017)
ImageNet; NOAA Nighttime Lights; Google Static Maps satellite imagery Proxy Task (Night-Light Prediction); Poverty Regression Survey correlation vs LSMS; MAE Xie et al. (2016)
Gold annotations from Amazon Mechanical Turk for English, Hindi, Italian, Portuguese; Wikipedia corpora Temporal Grounding Hour-range accuracy vs gold annotations Shwartz (2022)
FORK test set Commonsense QA (Culinary) Accuracy on US vs non-US probes; Statistical significance Palta and Rudinger (2023)
TV show dialogues; Cross-culture shows; LDC conversational corpora Norm Extraction; Self-Verification; Grounding AUC for grounding; human-judged best-norm selection Fung et al. (2024)
Reddit corpus of 61,981 users; Word association benchmarks Demographic-Conditioned LM; Word Association Perplexity; word-association accuracy Welch et al. (2020)
GeoMLAMA; FORK; CANDLE; DLAMA Cultural Commonsense QA; Country Prediction; Commonsense Verification Accuracy gaps; uniformity analysis Shen et al. (2024)
Concept and Application dataset Image Transcreation; Cultural Adaptation Human evaluation (relevance; meaning preservation) Khanuja et al. (2024)
WorldCuisines Multilingual VQA (Dish Prediction; Origin Prediction) Accuracy; adversarial context drop Winata et al. (2025)
CVQA Multilingual Visual QA Accuracy; answer-matching metrics; performance drop analysis Romero et al. (2024)
LAION, GeoDE, DollarStreet Annotation Suggestion; Data Selection Annotation cost vs quality; coverage of cultural features Ignat et al. (2024a)
Value-relevant outputs from 8 LLMs; Reference human value distributions from surveys Value Alignment Probing; Distribution Mapping Correlation with survey ground truth Cahyawijaya et al. (2025)
Table 8: Representative datasets and evaluation practices across cultural bias NLP tasks.
To curate a representative set of datasets and evaluation strategies for gender bias in NLP, we selected papers that span a wide spectrum of model architectures (from static embeddings to transformer-based LMs) and task settings (e.g., coreference resolution, occupation classification, translation, multi-agent interactions). The selection emphasizes both foundational work and recent advancements that shaped current methodologies. Early studies such as those by Bolukbasi et al. (2016) and Caliskan et al. (2017) were included for their role in establishing intrinsic bias probing techniques like WEAT. We also incorporated task-specific evaluations such as pronoun-drop metrics Stanovsky et al. (2019) and fairness gaps in classification tasks De-Arteaga et al. (2019). Recent papers were chosen to highlight emerging directions in LLM-based analysis, including causal interventions Cai et al. (2024), region-aware bias evaluations Borah et al. (2025), and multi-agent propagation frameworks Borah and Mihalcea (2024). Collectively, these works offer a diverse yet focused lens into how gender bias manifests and is measured in modern NLP systems.
We present a subset of representative datasets and evaluation benchmarks for cultural bias in Table 8.
To identify relevant datasets and benchmark efforts addressing the needs of people in underrepresented communities, like people with accessibility needs, we reviewed domain-specific surveys and high-impact papers (e.g., based on citation count and venue of publication). To focus on specific disability dataset, the literature was retrieved using search queries such as “NLP accessibility datasets”, “speech recognition for dysarthria”, “text simplification benchmark”, and “sign language translation dataset” in Google Scholar and ACL Anthology. When disability-specific datasets were not available, we included in the table the most widely used datasets for the corresponding NLP task. We present a representative collection of works in Table 9.
Datasets NLP Task Evaluation Metrics Reference
Newsela, WikiLarge,
WikiSmall, ASSET, MUSS,
SimplicityDA, Simple Wiki Text Simplification SARI, FKGL, BLEU, BERTScore, Human Ratings Al-Thanyyan and Azmi (2021)
LJSpeech, VCTK, LibriTTS, Blizzard Challenge, HiFi-TTS, M-AILABS, CSS10, AISHELL Text-to-Speech MOS (Mean Opinion Score), Intelligibility Score, Naturalness, Word Error Rate (WER), MCD (Mel-Cepstral Distortion) Khanam et al. (2022); Kumar et al. (2023)
EasyCall, UASpeech, TORGO, CSLU Dysarthric, DEED,
L2-ARCTIC, Google Project Euphonia Speech Recognition WER (Word Error Rate), CER (Character Error Rate), Accuracy, Intelligibility, Real-time Factor (RTF) Alharbi et al. (2021); Li et al. (2022)
DSBI, Smart Braille Converter Corpus, Tamil-Braille Dataset Braille Processing Accuracy, Precision, Recall, BLEU (for translation), OCR Error Rate Ali (2023)
MS COCO, VizWiz, Multi30K, Flickr8k, Flickr30k, STAIR Captions, TextCaps, OpenSubtitles Image Captioning and Subtitling BLEU, METEOR, ROUGE, CIDEr, SPICE, Human Ratings (fluency, adequacy), Caption Accuracy Stefanini et al. (2021); Ghandi et al. (2023)
VizWiz VQA, TDIUC, VQA-Med, OK-VQA, GQA, TextVQA, DocVQA, OViQA, PathVQA, Question Answering Accuracy, VQA Score, BLEU, ANLS (Average Normalized Levenshtein Similarity), Human Ratings Gurari et al. (2018); Chen et al. (2022); Huh et al. (2024)
PHOENIX14T, RWTH-PHOENIX-Weather, CSL-Daily, RWTH-BOSTON-104, ASLG-PC12, OpenASL, RWTH-SLT, Sign2Text, How2Sign, AUTSL Sign Language BLEU, ROUGE, METEOR, WER, Sign Error Rate (SER), Gloss Accuracy, Human Ratings Yin et al. (2021); Tao et al. (2024).
Table 9: Representative datasets and evaluation practices across accessibility-related NLP tasks.
A.8 Environmental Harms
Given the absence of a survey paper in this field, we began our review with research presented at the inaugural https://aclanthology.org/events/climatenlp-2024/. Additionally, we have included follow-up studies conducted by researchers within the ClimateNLP community. A significant portion of the work in this domain focuses on classifying climate-related claims, text, or stances. Detecting misinformation has emerged as a prominent topic, often tackled via external source verification and question-answering approaches. Information extraction, whether related to quantitative features or narrative insights, also plays a key role. The table 10 below highlights the key papers associated with these tasks.
Datasets NLP Task(s) Evaluation Metrics Reference
Global Stocktake Dataset from Climate Policy Radar Topic modelling Cosine similarity  Sietsma et al. (2023)
SumIPCC: topic-annotated summaries and relative paragraphs from climate change reports Aspect-based summarisation Mean Reciprocal Rank (MRR), Carburacy-reweighted ROUGE score  Ghinassi et al. (2024)
ClimateFever: real-world climate climate change claims with evidence sentences from Wikipedia Text classification Label-accuracy, F1, precision, recall  Diggelmann et al. (2020)
TCFD-category labeled sentences from firms’ annual reports, sustainability-, climate- or TCFD-reports and firms’ webpage Text classification Accuracy  Bingler et al. (2022)
CORP: paragraphs from common news, research articles and climate reporting of companies Language modelling, text classification, sentiment analysis, fact-checking Average cross-entropy loss, average validation loss, weighted F1 score  Leippold et al. (2022)
Reduction target claims collected by Net Zero Tracker Text classification Accuracy, precision, recall, F1 score  Schimanski et al. (2023)
HLEG reports, Net Zero Stocktake reports, Corporate Climate Responsibility Monitor Reports from NewClimate Institute Q&A, text generation Expert evaluation scores for quality, factual accuracy, relevance Hsu et al. (2024)
Corpus created from curated sources compiled by ERASMUS.AI (Pretraining), collection of scientific reports and papers (RAG), ClimaBench (Downstream tasks): collection of climate-related datasets for classification Language modelling, Q&A, various downstream tasks (classification etc.) Cross-entropy loss, average validation loss, weighted F1-score, BLEU scores, human evaluation  Thulke et al. (2024)
Sixth Assessment Reports (AR6) of IPCC Q&A, text generation Accuracy score given by experts  Vaghefi et al. (2023)
Text from IPCC, WMO, AbsCC (climate change abstracts), 1000S (abstracts by top 1000 climate scientists) Text classification Averaged micro-F1 score with different classification levels Leippold et al. (2025)
SemTabNet: tables from over 10K corporate ESG reports obtained using Deep Search toolkit Information extraction Tree Similarity Score  Mishra et al. (2024)
Corporate annual and sustainability reports Quantitative information extraction Custom report-level metrics evaluating retrieval and accuracy of extractions  Dimmelmeier et al. (2024)
ClimateQA: climate claims Q&A Manual inspection, ROUGE-L recall score, conditional probability, truth ratio, GPT-Match, GPT-Contradiction, AlignScore Fore et al. (2024)
Climate change-related questions from Google Trends, Skeptical Science, synthetic questions from English Wikipedia Evaluating LLMs’ responses Presentational (style, clarity, correctness, tone) and epistemological (accuracy, specificity, completeness and uncertainty) properties  Bulian et al. (2024)
Corporate sustainability reports Text summarization, Text scoring, Q&A Human evaluation for hallucinations, ROUGE precision score  Ni et al. (2023)
Climate-related questions (question-source-answer pairs), sustainability report dataset (report-paragraph-question pairs) Information extraction Recall@K, Precision@K, F1@K for different top K values Schimanski et al. (2024)
Dataset of corporate climate policy engagement documents collected by LobbyMap Information extraction (query, stance and evidence page indices), classification Strict F-score, page overlap F-score, document F-score, Morio and Manning (2023)
Facebook ads related to climate change (oil and gas sector), text and spend, impressions, demographic and regional distribution Multi-label classification Overall and sub-category specific F-score Holder et al. (2023); Rowlands et al. (2024)
News articles in English and Mandarin Information extraction (narrative features) Human evaluation, ROUGE-1, ROUGE-L, cosine similarity Zhou et al. (2024)
Table 10: Selection of influential datasets and papers in the ClimateNLP domain, ranging from topic modelling over text classification to question answering and information extraction
Appendix B Global Goals
We include an overview of the Sustainable Development Goals (SDGs) in Figure 2, which apply to many of the NLP applications discussed in this paper.
Figure 2: Overview of the SDG goals. Source: https://sdgs.un.org/goals
Appendix C Global Risks
We present the key global risks categorized by domain, as outlined in the Global Risks Report 2025 by the World Economic Forum. [^7] https://www.weforum.org/publications/global-risks-report-2025 Each domain-specific table in Figure 3 highlights major threats along with their definitions. We also provide the global risks ranked by severity over the short and long term in Figure 4.
Figure 3: Domain-specific global risks according to the Global Risks Report 2025. Each table illustrates key risks in societal, technological, geopolitical, environmental, and economic domains respectively.
Figure 4: Global risks ranked by severity over the short and long term. Reproduction of Figure FIGURE C from: the Global Risks Report 2025
Appendix D Guidelines for authors
This work involved contributions from many collaborators with diverse backgrounds, who surveyed various topics and guided the writing process. After extensive discussions in the project’s early phase, we developed the author guidelines shown in Figure 5. We share these here for transparency reasons and to assist other researchers undertaking similar multidisciplinary efforts.
Project Guidelines Your task is to write a section (or subsection) based on a topic of your expertise. To ensure consistency across all sections, we will first collaboratively collect key papers, identify the needs they address, and outline the NLP methodologies they apply. This information, along with our notes, will be organised in a shared spreadsheet. Once the research is mapped out, we will proceed to draft the sections, following a consistent structure that highlights both current opportunities and existing challenges in the field. Please work on your chosen topic tab and fill in the shared [link]spreadsheet. Below are the main suggestions to help guide your input: Spreadsheet Columns: Main Field Papers – Select key papers in the area. We recommend starting from potential existing surveys in the field to identify key papers. If no surveys exist, we recommend using a keyword search and manual filtering of the most influential papers using the number of citations if needed. Social Needs Covered – Societal challenges addressed. We need this information for Section 2. So, please update the column in your tab accordingly, even if you don’t use it in your section. Popular Datasets – Commonly used datasets NLP Task(s) – How tasks are defined (generation, classification, etc.) NLP Methodology (Existing) – Methods used in literature Evaluation – Evaluation setup and metrics Limitations – Limitations of current work Challenges / Open Questions – Remaining gaps or issues Expected NLP Impact / Suggestions – Potential contributions and ideas NLP Methodology Potential – Methodological insights or improvements Suggested structure for your section: Methodology: – Dataset, approach, evaluation Limitations: – Challenges, Critical analysis, and Future Work Opportunities: – Suggestions, Open Questions, Impact, Broader relevance and Impact in NLP Suggested length for your topic-section: We aim for three main discussions per topic: methodology, limitations, and opportunities. Please use a table for the methodology section that will be added in the appendix. We can aim for a maximum of 4 paragraphs for your chosen topic: general intro, methodology paragraph (going to the appendix), challenges, opportunities. In the main paper, you have one column each. Keep in mind that we might keep a shorter version in the final revisions.
Figure 5: Guidelines for the authors of the paper. Please reach out for any clarification.

### v2

@@
-NLP for Social Good: A Survey of Challenges, Opportunities, and Responsible Deployment
+NLP for Social Good: A Survey and Outlook of Challenges, Opportunities, and Responsible Deployment
 Abstract
-Recent advancements in large language models (LLMs) have unlocked unprecedented possibilities across a range of applications.
-However, as a community, we believe that the field of Natural Language Processing (NLP) has a growing need to approach deployment with greater intentionality and responsibility.
-In alignment with the broader vision of AI for Social Good Tomašev et al.
-(2020), this paper examines the role of NLP in addressing pressing societal challenges.
-Through a cross-disciplinary analysis of social goals and emerging risks, we highlight promising research directions and outline challenges that must be addressed to ensure responsible and equitable progress in NLP4SG research.
-NLP for Social Good: A Survey of
+Natural language processing (NLP) now shapes many aspects of our world, yet its potential for positive social impact is underexplored.
+This paper surveys work in “NLP for Social Good" (NLP4SG) across nine domains relevant to global development and risk agendas, summarizing principal tasks and challenges.
+We analyze ACL Anthology trends, finding that inclusion and AI harms attract the most research, while domains such as poverty, peacebuilding, and environmental protection remain underexplored.
+Guided by our review, we outline opportunities for responsible and equitable NLP and conclude with a call for cross-disciplinary partnerships and human-centered approaches to ensure that future NLP technologies advance the public good.
+NLP for Social Good: A Survey and Outlook of
 Challenges, Opportunities, and Responsible Deployment
-Antonia Karamolegkou1, Angana Borah2, Eunjung Cho3, Sagnik Ray Choudhury4, Martina Galletti5,6, Rajarshi Ghosh7, Ishani Mondal8, Pranav Gupta9, Oana Ignat10, Priyanka Kargupta11, Neema Kotonya12, Hemank Lamba12, Sun-Joo Lee13, Arushi Mangla10, Deniz Nazarova14, Poli Nemkova4,13, Dina Pisarevskaya15, Naquee Rizwan16, Nazanin Sabri17, Dominik Stammbach18, Anna Steinberg19,20, David Tomás21, Steven R Wilson22, Bowen Yi2,23, Jessica Zhu8, Arkaitz Zubiaga15, Anders Søgaard1, Alexander Fraser20,27, Zhijing Jin3,24,25,26, Rada Mihalcea2, Joel R.
-Tetreault12, Daryna Dementieva27 1University of Copenhagen 2University of Michigan-Ann Arbor 3ETH Zurich 4 University of North Texas 5Sony Computer Science Laboratories - Paris 6University of Rome “La Sapienza” 7Algoverse 8University of Maryland, College Park 9Lowe’s 10Santa Clara University 11University of Illinois Urbana-Champaign 12Dataminr 13 United Nations Development Programme (UNDP) 14University of Washington 15Queen Mary University of London 16IIT Kharagpur 17University of California San Diego 18 Princeton University 19LMU Munich 20Munich Center for Machine Learning (MCML) 21University of Alicante 22University of Michigan-Flint 23University of Southern California 24Max Planck Institute for Intelligent Systems, Tübingen 25Vector Institute 26University of Toronto 27Technical University of Munich Correspondence: antka@di.ku.dk, daryna.dementieva@tum.de
+Antonia Karamolegkou1, Angana Borah2, Eunjung Cho3, Sagnik Ray Choudhury4, Martina Galletti5,6, Pranav Gupta7, Oana Ignat8, Priyanka Kargupta9, Neema Kotonya10, Hemank Lamba10, Sun-Joo Lee11, Arushi Mangla8, Ishani Mondal12, Fatima Zahra Moudakir13,14,15, Deniz Nazar16, Poli Nemkova4,11, Dina Pisarevskaya17, Naquee Rizwan18, Nazanin Sabri19, Keenan Samway13, Dominik Stammbach20, Anna Steinberg Schulten21,22, David Tomás23, Steven R Wilson24, Bowen Yi2,25, Jessica Zhu12, Arkaitz Zubiaga17, Anders Søgaard1, Alexander Fraser22,26, Zhijing Jin3,13,14,15, Rada Mihalcea2, Joel R.
+Tetreault10, Daryna Dementieva22,26 1University of Copenhagen 2University of Michigan-Ann Arbor 3ETH Zurich 4University of North Texas 5Sony Computer Science Laboratories - Paris 6University of Rome “La Sapienza” 7Lowe’s 8Santa Clara University 9University of Illinois Urbana-Champaign 10Dataminr 11United Nations Development Programme (UNDP) 12University of Maryland, College Park 13Max Planck Institute for Intelligent Systems, Tübingen 14Vector Institute 15University of Toronto 16University of Washington 17Queen Mary University of London 18IIT Kharagpur 19University of California San Diego 20Princeton University 21LMU Munich 22Munich Center for Machine Learning (MCML) 23University of Alicante 24University of Michigan-Flint 25University of Southern California 26Technical University of Munich Accepted to EACL 2026 (Main Conference), Rabat, Morocco.
+Correspondence: antoniakrm16@gmail.com, daryna.dementieva@tum.de
 1 Introduction
@@
 To fully realize the potential of NLP, it is essential to look beyond technical achievements and reframe tasks around pressing societal needs.
-We draw on insights from the United Nations Sustainable Development Goals [^1] https://sdgs.un.org/goals (UN SDGs) and the 2025 Global Economic Risks Report [^2] https://www.weforum.org/publications/global-risks-report-2025/digest (GR) to provide a foundation for an interdisciplinary recontextualization of NLP, encouraging reflection on how language technologies intersect with today’s most pressing challenges.
+We draw on insights from the United Nations Sustainable Development Goals [^1] https://sdgs.un.org/goals (SDGs) and the 2025 Global Risks (GRs) Economic Report [^2] https://www.weforum.org/publications/global-risks-report-2025/digest to provide a foundation for an interdisciplinary recontextualization of NLP, encouraging reflection on how language technologies intersect with today’s most pressing challenges.
 We selected these two agendas as, from a social good perspective, UN SDGs offer a global framework for fostering peace and prosperity for people and the planet.
 However, while highly influential, these goals were established in 2015—prior to the rapid advancements in artificial intelligence.
-To contextualize them within today’s technological landscape, we also draw on insights from the 2025 GR Report, which highlights both the transformative potential and the emerging global risks (GRs) associated with technology and information processing.
+To contextualize them within today’s technological landscape, we also draw on insights from the 2025 GR Report, which highlights both the transformative potential and the emerging global risks associated with technology and information processing.
+The resulting mapping of NLP application domains to SDGs and GRs is shown in Figure 1.
 Our effort builds on prior research that assesses the role of NLP through positive impact Hovy and Spruit (2016); Jin et al.
@@
 (2025).
-Thus, our research goal in this work is threefold:
-  1. RQ1:
-    What NLP-based solutions already support positive social impact?
-  2. RQ2:
-    What are the main challenges in developing NLP for positive social impact?
-  3. RQ3:
-    What promising directions are currently overlooked?
-To address these questions, we identify nine key NLP research directions, as illustrated in Figure 1, using the relevant SDGs and GRs.
-The following sections summarize research tasks, [^3] Details about methodology and motivation of paper selection can be found in the Appendix A.
-the open challenges and opportunities within each domain, concluding with overarching open research directions and recommended community actions for proactive NLP4SG efforts.
+Thus, our research goal in this work is threefold: RQ1—what NLP-based solutions already support positive social impact, RQ2—what challenges arise in developing them, and RQ3—what promising directions remain overlooked?
+We first analyze publication trends in the ACL Anthology across nine NLP research directions (Figure 2), revealing uneven growth across domains.
+For each domain, we review existing work and current challenges, [^3] Further statistics and details about survey methodology and paper selection are provided in Appendix B.
+and present an outlook on opportunities that reflects our perspective on how the field can advance.
+We conclude by synthesizing overarching research directions and community actions to encourage more proactive NLP4SG efforts.
+Figure 2: Number of ACL Anthology papers per domain showing publication volume and growth trajectories over time (2019–2025).
 2 [Uncaptioned image]  Healthcare
-NLP in healthcare is a space that targets problems associated with SDG3: Good Health and Well-being by improving healthcare delivery and outcomes.
-It also targets critical global risks (GR4), including the decline in health and well-being, health workforce shortages, and infectious disease outbreaks.
-Mental Health
-Given that the need for mental health support increasingly remains unmet, NLP systems, specifically LLMs, address an access gap for users who are limited in time and resources Tong et al.
+NLP can help address challenges linked to SDG3 (Good Health and Well-being) by improving healthcare delivery and outcomes; and to GR4 (Societal Risks), including health decline, health workforce shortages, and infectious disease outbreaks.
+Mental Health.
+As healthcare access remains uneven, NLP systems, particularly LLMs, offer scalable means to reduce support gaps for those limited in time and resources Tong et al.
 (2023); Hua et al.
@@
 (2024); Liu et al.
-(2025a).
-Physical Health
+(2025b).
+Physical Health.
 Prior work has focused on physical well being, using social media mining for tracking physical activity, sleep patterns Sakib et al.
@@
 Challenges:
-Key challenges include health-related data, which is often scarce, sensitive, and affected by systemic biases, with language limitations and underrepresentation of marginalized groups raising ethical and privacy concerns (Ford et al., 2019; Shakeri Hossein Abad et al., 2022; Gunal et al., 2025); evaluation, which must go beyond accuracy to reflect fairness, contextual understanding, human-centered values (e.g.
+Key challenges involve the availability of health-related data, which is scarce, sensitive, and often biased, with limited language coverage and marginalized group representation raising ethical and privacy concerns (Ford et al., 2019; Shakeri Hossein Abad et al., 2022; Gunal et al., 2025); the evaluation framework, which must go beyond accuracy to reflect fairness, contextual understanding, human-centered values (e.g.
 empathy), and ensure reproducibility; and long-term user impact, as reliance on LLMs for sensitive tasks risks sycophancy, ELIZA effects Ekbia (2008), and overdependence.
@@
 (2022); Brown and Halpern (2021).
+There is growing evidence that LLMs lack the stability, contextual grounding, and ethical guarantees required for autonomous clinical or therapeutic decision-making (Zhao et al., 2024; Iftikhar et al., 2025; Cui et al., 2025), highlighting the need for informed, cautious, and supervised use of such tools.
 Opportunities:
-Future work should extend beyond text to multimodal approaches that integrate speech prosody, facial expressions, physiological signals, and other data sources like sensors, wearables, or images for richer context and personalization Puce et al.
+Future work should continue to emphasize multimodal approaches that jointly leverage text, speech prosody, facial expressions, physiological signals, and other data sources like sensors, wearables, or images for richer context and personalization Puce et al.
 (2025).
@@
 (2021).
-Crucially, interdisciplinary research is needed to build AI-augmented therapeutic frameworks that complement human care—especially for underserved communities.
+Importantly, interdisciplinary research is needed to build AI-augmented therapeutic frameworks that complement human care—especially for underserved communities.
 3 [Uncaptioned image]  Education
-The integration of LLMs into education offers tools that support SDG4: Quality Education, and also addresses a variety of risks by fostering informed and critical individuals (GR4, GR5).
+The integration of LLMs into education offers tools that support SDG4: Quality Education, and also addresses a variety of societal risks by fostering informed and critical individuals (GR4).
 NLP is used for automated feedback Jurenka et al.
@@
 (2022).
-AI Literacy
 As AI systems become more embedded in everyday life, AI literacy is essential.
 A recent review (Yang et al., 2025) draws connections with digital, data, and algorithmic literacies.
-However, NLP-driven efforts for AI literacy remain scattered (Long and Magerko, 2020; Congress, 2023; Moorkens et al., 2024; Tapo et al., 2025; Korea Education and Research Information Service, 2025)
+However, NLP-driven efforts for AI literacy remain scattered (Long and Magerko, 2020; Congress, 2023; Moorkens et al., 2024; Tapo et al., 2025; Korea Education and Research Information Service, 2025).
 Challenges:
@@
 (2024); Kargupta et al.
-(2024)—hinder the effective integration of NLP into education.
+(2024) hinder the effective integration of NLP into education.
 Mixed perceptions and mistrust toward AI also remain a barrier Nader et al.
@@
 Opportunities:
-Further research should be encouraged for specific groups, such as teachers Du et al.
-(2024), students Shen and Cui (2024), and other professionals Lo (2024).
-Another important opportunity is the alignment with expert-annotated pedagogical traces and grounding evaluations in curriculum-linked outcomes, to bridge the gap between subject expertise and pedagogical effectiveness Macina et al.
+We believe LLMs can be a useful educational tool under sustained human oversight, but it cannot replace the relational, pedagogical, and contextual roles of human educators (Martynova et al., 2025; Calvert et al., 2025).
+Future work should focus on specific groups such as teachers Du et al.
+(2024), students Shen and Cui (2024), and other professionals Lo (2024) to obtain different perspectives on education.
+Moreover, aligning with expert-annotated pedagogical traces and grounding evaluations in curriculum outcomes to bridge subject expertise and pedagogical effectiveness Macina et al.
 (2025); Lucy et al.
 (2024).
-Human-in-the-loop approaches have also shown promise in scaling expert strategies while preserving teacher agency, particularly benefiting underserved communities Wang et al.
+Human-in-the-loop methods can scale expert strategies while preserving teacher agency, particularly in underserved settings Wang et al.
 (2024c).
-Multi-agent simulations offer privacy-preserving environments to test classroom policies and assess equity metrics before real-world deployment Zhang et al.
+Multi-agent simulations provide privacy-preserving environments to test classroom policies and equity before real-world deployment Zhang et al.
+(2025), for instance, through Socratic planning agents that promote critical thinking over rote learning Kargupta et al.
 (2024).
-For example, Socratic planning agents encourage critical thinking over rote memorization, fostering deeper learning experiences Kargupta et al.
-(2024).
-Finally, community-driven initiatives are crucial for consolidating best practices for the development of safe and accountable educational AI systems Chu et al.
+Finally, community-driven efforts are key to consolidate best practices for safe and accountable educational AI Chu et al.
 (2025); Wen et al.
@@
 4 [Uncaptioned image]  Poverty
-Economic downturn (GR3) is the sixth highest-ranked GRs, with serious implications for poverty—one of the world’s most pressing challenges (Lister, 2021) and the focus of SDG1.
+Economic downturn (GR3) is the sixth highest-ranked GRs, with serious implications for poverty (SDG1)—one of the world’s most pressing challenges (Lister, 2021).
 Nearly 700 million people continue to live on less than \$2.15 per day (Hasell et al., 2024; World Bank Group, 2024).
@@
 Opportunities:
-To enable large-scale, global analysis of poverty, high-quality datasets labeled with income, socioeconomic status, or poverty indicators are essential, potentially gathered via data donations, user surveys, or social media statistics.
+To enable a global analysis of poverty, datasets labeled with income, socioeconomic status, or poverty indicators are essential, potentially gathered via data donations, user surveys, or social media statistics.
 [^4] https://datareportal.com/social-media-users Additionally, use-case-specific NLP applications can advance poverty research (Adauto et al., 2023).
@@
 5 [Uncaptioned image]  Peacebuilding
-Peacebuilding is essential to achieving SDG16, and despite escalating threats like state-based armed conflict (R1); NLP’s application in key tools for peacebuilding like conflict prediction, human rights monitoring, and physical safety remains limited and underexplored.
-Human Rights Violations
-NLP in human rights mainly involves classification and detection of violations across languages, such as Arabic Twitter (Alhelbawy et al., 2016), Jordanian Arabic social media (Khalafat et al., 2021), Russian and Ukrainian Telegram (Nemkova et al., 2023), and English Twitter (Pilankar et al., 2022).
-Other tasks include threat detection for human rights defenders (Ran et al., 2023) and interpretable forced labor identification (Guzman et al., 2024).
-LLMs show potential for zero- and few-shot annotation but require human oversight (Nemkova et al., 2025b).
-Conflict Prediction
-NLP in conflict prediction mainly involves extracting and forecasting conflict dynamics using structured datasets like UCDP and ACLED (Hegre et al., 2019).
-Tasks include topic modeling for early signals (Mueller et al., 2024), modeling conflict stages with LLMs (Croicu and von der Maase, 2025; Nemkova et al., 2025a), automated report generation via retrieval-augmented generation (Nemkova et al., 2025c), and conflict content extraction and classification with domain-adapted models like ConfliBERT (Hu et al., 2022).
-Physical Safety
-NLP systems support understanding and intervention in threats to physical safety.
-They can be used to analyze social media, police reports, surveillance footage, health records, and news articles.
-Techniques like topic modeling More and Francis (2021), sentiment analysis Blevins et al.
+Peacebuilding is essential to achieving SDG16, and despite escalating geopolitical threats (GR1), NLP tools for peacebuilding like human rights monitoring, conflict prediction, and physical safety remains limited and underexplored.
+Human Rights Violations.
+NLP for human rights focuses on detecting violations across languages and platforms, from Arabic and Jordanian social media (Alhelbawy et al., 2016; Khalafat et al., 2021) to Russian and Ukrainian Telegram (Nemkova et al., 2023) and English Twitter (Pilankar et al., 2022).
+Beyond general classification, efforts target threats to defenders (Ran et al., 2023a), forced labor with interpretable models (Guzman et al., 2024b), human trafficking (Liu et al., 2023; Saxena et al., 2023, 2024), dehumanization (Caporusso et al., 2024; Burovova and Romanyshyn, 2024), and other abuses (Yu, 2022; Guzman et al., 2024a; Sorato et al., 2024; De Paula et al., 2024; Wang, 2024).
+Conflict Prediction.
+NLP supports the extraction and forecasting of conflict dynamics (Halkia et al., 2020; Stoehr et al., 2021; Alsarra et al., 2023; Sathvik et al., 2024).
+Methods span from topic modeling for early warning (Mueller and Rauh, 2018; Mueller et al., 2024a, b) to LLM-based approaches modeling escalation and resolution, including emerging agentic paradigms (Croicu and von der Maase, 2025; Nemkova et al., 2025a; Nemkova and Albert, ).
+Studies also increasingly analyze civil unrest and protest dynamics (Sech et al., 2020; Chinta et al., 2021; Scharf et al., 2021; Raj et al., 2022; Siskou et al., 2022; Wiedemann et al., 2022; Loerakker et al., 2024; Olsen et al., 2024).
+Advances include domain-adapted encoders like ConfliBERT for improved extraction and classification (Hu et al., 2022) and automated situation reports via retrieval-augmentation (RAG) (Nemkova et al., 2025c).
+Recent work extends to diplomatic negotiation and argument mining, analyzing discourse in peace talks and diplomacy (Glaser et al., 2022; Rødven-Eide et al., 2023; Zaczynska and Stede, 2024; Zaczynska et al., 2024; Anisimova and Zikánová, 2024; Poiaganova and Stede, 2025) by modeling pathways to resolution.
+Physical Safety.
+They have been used to analyze relevant social media Al-Garadi et al.
+(2022a); Levy et al.
+(2022); Blandfort et al.
+(2019); Blevins et al.
+(2016), police reports Karystianis et al.
+(2019), surveillance footage Kumari et al.
+(2023), health records Borger et al.
+(2022); Botelle et al.
+(2022); MacPhaul et al.
+(2023), reporting systems Chew et al.
+(2023); Arseniev-Koehler et al.
+(2022), and news articles Pavlick et al.
+(2016).
+Techniques such as topic modeling More and Francis (2021), sentiment analysis Blevins et al.
 (2016), entity extraction Pavlick et al.
@@
 Challenges:
-First, Data and Domain Complexity: references to violations and threats are often indirect or euphemistic, complicating detection.
-Labeled multilingual datasets—especially for low-resource or sensitive regions—are scarce ALSaif and Alotaibi (2019); MacPhaul et al.
-(2023); Parker (2020); Chang et al.
-(2018), and causes are often multifaceted and extend beyond textual signals.
-Annotation is labor-intensive with expert disagreement common Blandfort et al.
+Peacebuilding poses several challenges.
+Language and Context: threats and violations are often euphemistic, coded, or strategically disguised, especially in authoritarian or conflict settings Nemkova et al.
+(2023); Ran et al.
+(2023b); Nemkova et al.
+(2025b).
+Data is fragmented across low-resource languages and dialects, with annotation hindered by political sensitivity and expert disagreement Blandfort et al.
 (2019); Levy et al.
@@
 (2022); Chew et al.
-(2023), while evaluation suffers from the absence of standardized benchmarks and reliable ground truth, limiting model validation and comparison.
-Second, Model Limitations: NLP models struggle with code-switching, generalization across languages, regions, and out-of-distribution data, and often reflect training biases.
-The black-box nature of LLMs reduces transparency, hindering their deployment in high-stakes contexts, with real-world application still in early stages and lacking validated best practices.
-Lastly, Ethical and Evaluation Challenges: misclassifications can lead to legal or advocacy risks, highlighting the need for rigorous evaluation and ethical care.
+(2023).
+Temporal Volatility: conflicts evolve rapidly, as protests, ceasefires, or escalations outpace model adaptation, leaving systems trained on historical data prone to domain drift ALSaif and Alotaibi (2019); MacPhaul et al.
+(2023); Parker (2020); Chang et al.
+(2018).
+High-Stakes Ethics: misclassifications may expose activists, mislead humanitarian response, or legitimize violence, demanding stricter evaluation and oversight than other domains Kumari et al.
+(2023); Al-Garadi et al.
+(2022a); Chang et al.
+(2018).
+Since evidence for robust generalization remains limited, cautious interpretation and sustained human expertise are essential in deployment.
 Opportunities:
-On a computational level, tools such as conflict prediction models and systems for detecting human rights violations can support evidence-based policymaking, inform the development and targeting of interventions, and enable rapid programmatic adjustments in response to evolving conditions.
-Developing multilingual and cross-regional models will enable more comprehensive monitoring.
-Further, NLP methods can contribute to operational efficiency for implementing organizations.
-For example, LLMs offer capabilities for rapid document review and synthesis, facilitating timely and contextually informed analysis.
-RAG systems can further enhance information access by efficiently extracting relevant content from document repositories.
-Additionally, LLMs can support scenario generation, situational analysis, and strategic planning.
-Lastly, a dedicated peacebuilding workshop could offer a crucial forum for interdisciplinary collaboration on NLP tools for humanitarian aid and crisis response.
+We believe that models such as conflict predictors and human rights violation detectors can support policymaking, interventions, and programmatic adjustments across multilingual and cross-regional contexts ALSaif and Alotaibi (2019).
+NLP methods can also enhance operational efficiency: LLMs enable rapid document review and synthesis for timely, context-aware decisions Borger et al.
+(2022); Chew et al.
+(2023), while RAG systems improve access to relevant information Nemkova et al.
+(2025c).
+LLMs further assist with scenario generation, situational analysis, and strategic planning Nemkova et al.
+(2025c).
+Finally, a dedicated peacebuilding NLP workshop could provide a crucial interdisciplinary forum for developing tools for humanitarian aid and crisis response.
 6 [Uncaptioned image] Environment Protection
-NLP offers scalable tools for climate mitigation and adaptation—supporting key SDGs (SDG6, SDG11, SDG12, SDG14, SDG15) and addressing critical global environmental risks (R2)—by extracting insights from unstructured text like scientific papers, policy reports, and assessments (on Climate Change, 2022).
+NLP offers scalable tools for climate mitigation and adaptation—supporting key SDGs (SDG6, SDG7, SDG11, SDG12, SDG14, SDG15) and addressing critical global environmental risks (R2)—by extracting insights from unstructured text like scientific papers, policy reports, and assessments (on Climate Change, 2022).
 Techniques such as topic modeling (Sietsma et al., 2023), summarization (Ghinassi et al., 2024), and classification (Varini et al., 2020; Stammbach et al., 2023; Bingler et al., 2022; Schimanski et al., 2023) enable analysis of diverse datasets.
@@
 Extracting quantitative information, particularly from sustainability reports, remains challenging due to the complex, multi-modal and non-standardized nature of the data sources, leading to pipelines specifically designed for information extraction from tables (Mishra et al., 2024; Dimmelmeier et al., 2024).
-Furthermore, LLMs are prone to hallucination (Vaghefi et al., 2023) or are contaminated with false, conflicting or outdated climate information (Fore et al., 2024)Bulian et al.
-(2024) propose an evaluation framework based on presentational and epistemological qualities to evaluate quality and factual accuracy of LLMs.
+Furthermore, LLMs, prone to hallucination (Vaghefi et al., 2023), often output false, conflicting or outdated climate information (Fore et al., 2024)Bulian et al.
+(2024) evaluate the quality and factual accuracy of LLMs on climate information using a framework based on presentational and epistemological qualities.
 Opportunities:
-NLP is enabling cross-disciplinary collaboration, bridging domain experts with the fields of computer science, social science, and economics to analyze climate policies (Gandhi et al., 2024) or bringing the scientific community and non-governmental organization together to uncover narratives in public climate discourse (Gehring and Grigoletto, 2023; Rowlands et al., 2024).
+NLP can enable cross-disciplinary collaboration, bridging domain experts from fields such as computer science, social science, and economics to analyze climate policies (Gandhi et al., 2024) or bringing the scientific community and non-governmental organizations to uncover narratives in public climate discourse (Gehring and Grigoletto, 2023; Rowlands et al., 2024).
 This is particularly important in expanding research to include cross-cultural and multilingual perspectives (Zhou et al., 2024; Bird et al., 2024).
+Lastly, NLP can inform other fields through climate reporting (Hershcovich et al., 2022b) and support sustainable behavior change (Chockkalingam et al., 2025).
 7 [Uncaptioned image] Inclusion and Inequalities
-Inclusive NLP that accounts for demographic, linguistic, and accessibility diversity is key to reducing systemic inequalities.
-These efforts support SDG5 and SDG10, and address inequality—a top global concern ranked among the most severe short- and long-term risks (GR4).
-Gender Bias
-Gender bias in NLP—especially in hiring De-Arteaga et al.
-(2019), healthcare, and moderation—can reinforce stereotypes and harm marginalized groups.
-Initial work focused on static embeddings using gender direction Bolukbasi et al.
-(2016) and WEAT Caliskan et al.
-(2017), later extending to contextual tasks like coreference Zhao et al.
-(2018) and occupation classification De-Arteaga et al.
+Inclusive and equitable language technologies are central to addressing systemic inequalities (disparities in how individuals or groups are represented, served, or affected), often reflecting socioeconomic, cultural, linguistic, and accessibility hierarchies.
+These efforts align with key sustainable development goals (SDG5, SDG10) and address inequality, ranked among the most severe societal risks in both the short and long term (GR4).
+A growing body of survey work examines bias and fairness in NLP, cataloguing demographic bias (gender, race, ethnicity, age, sexual orientation, disability, and socioeconomic status) (Gupta et al., 2024), its origins (Brunet et al., 2019; Hovy and Prabhumoye, 2021), alongside detection, quantification and mitigation methods  Stanczak and Augenstein (2021); Devinney et al.
+(2022); Bartl et al.
+(2025); Gallegos et al.
 (2024).
-Recent studies examine LLM-agent interactions Borah and Mihalcea (2024) to track how biases spread.
-Mitigation includes post-processing (e.g., fine-tuning, DPO), interpretability, causal inference Attanasio et al.
+Studies have focused on bias in model representations Bolukbasi et al.
+(2016); Caliskan et al.
+(2017), creating bias detection benchmarks (Zhao et al., 2018; Parrish et al., 2022), and testing models in high-stakes settings De-Arteaga et al.
+(2019); Cross et al.
+(2024).
+More recent research addresses bias in agentic systems Borah and Mihalcea (2024), proposing mitigation strategies including post-processing, interpretability-driven and causal approaches Attanasio et al.
 (2023); Cai et al.
-(2024), and counterfactual or intervention-based prompting Plyler and Chi (2025).
-Bias is commonly assessed via fairness gaps (e.g., F1, error disparities) or associations (e.g., WEAT, Bias-Score).
-Cultural and Linguistic Representation
-NLP systems often fail to capture the full diversity of human cultures and languages.
-Culture’s complexity - including evolving norms, values, and worldviews - varies widely across communities (Saha et al., 2025).
-Despite culture’s crucial role in communication, AI models frequently reflect incomplete cultural perspectives, overrepresenting dominant or cultural linguistic narratives typically from high-resource languages or communities (Hershcovich et al., 2022a; Mihalcea et al., 2025).
-Underrepresented Communities
-Users with diverse accessibility needs remain underserved, continuing to face significant communication barriers Khanuja et al.
+(2024), as well as counterfactual prompting Plyler and Chi (2025).
+Beyond demographic bias, inequalities also arise from the limited support of underrepresented communities with diverse accessibility needs Khanuja et al.
 (2023a).
-NLP has been used in a wide range of assistive applications, including as augmentative communication Park et al.
+NLP has been applied to a wide range of assistive technologies, including augmentative communication Park et al.
 (2022), text simplification Espinosa-Zaragoza et al.
-(2023), text-to-speech Kumar et al.
-(2023), speech recognition Li et al.
+(2023), text-to-speech and speech recognition Kumar et al.
+(2023); Li et al.
 (2022), braille processing Tejesh et al.
@@
 Challenges:
-Many issues revolve around binary gender assumptions and limited cross-linguistic/cultural generalizability Stanovsky et al.
+Many challenges stem from simplifying assumptions, including binary gender representations, coarse socioeconomic proxies (Bassignana et al., 2025), and unequal access to digital resources Narayanan Venkit (2023).
+Limited cross-linguistic and cross-cultural generalizability further constrains model reliability Stanovsky et al.
 (2019); Adilazuarda (2024).
-Western-centric resources and lack of intersectional frameworks reinforce biases and marginalization  Kleinberg and Raghavan (2021); Sewunetie et al.
+Culture—encompassing evolving norms, values, and worldviews—varies widely across communities (Saha et al., 2025), yet NLP systems often reflect incomplete cultural perspectives, overrepresenting dominant linguistic narratives from high-resource languages and regions (Hershcovich et al., 2022a; Karamolegkou et al., 2024; Mihalcea et al., 2025).
+Western-centric resources and the lack of intersectional frameworks reinforce marginalization Kleinberg and Raghavan (2021); Sewunetie et al.
+(2024), while compounded forms of discrimination—arising at the intersection of race, gender, social status, and disability—remain underexplored Stanczak and Augenstein (2021); Guo and Caliskan (2021); Wald (2021).
+Progress is further hindered by scarce multimodal datasets, particularly for Braille and sign languages Hutchinson and Prabhakaran (2020); De Sisto et al.
+(2022); Karamolegkou et al.
+(2025b), limited interdisciplinary collaboration Kusters et al.
+(2020), and reliance on synthetic or prompt-generated data that obscures real-world biases Venkit et al.
+(2025); Morales et al.
 (2024).
-Implicit biases and compounded discrimination, e.g., at the intersection of race, gender and disability, are underexplored Stanczak and Augenstein (2021); Guo and Caliskan (2021); Wald (2021).
-Data challenges persist due to scarce multimodal datasets, especially for Braille and sign language Hutchinson and Prabhakaran (2020); De Sisto et al.
-(2022).
-Limited interdisciplinary collaboration slows progress Kusters et al.
-(2020).
-Reliance on synthetic or prompt-generated data hampers real-world validity and hides systemic biases Venkit et al.
-(2025); Morales et al.
-(2024), while opaque data sourcing blocks accountability Bender and Friedman (2018).
+Finally, opaque data sourcing practices limit accountability and reproducibility Bender and Friedman (2018).
 Opportunities:
@@
 (2023); Lauscher et al.
-(2022) and collaborative data collection with cultural experts and underrepresented groups Bird and Yibarbuk (2024); Newman-Griffis et al.
-(2024); UNICEF (2020); Hirmer et al.
+(2022) and collaborative data collection with cultural experts and underrepresented groups UNICEF (2020); Hirmer et al.
+(2021); Bird and Yibarbuk (2024); Newman-Griffis et al.
+(2024); Karamolegkou et al.
 (2024).
@@
 (2024d).
-8 [Uncaptioned image]  Online Harms
+8 [Uncaptioned image]  Digital Violence
 The recent ease of access to digital devices (like smartphones and those based on IoT) has fueled the spread of digital violence globally (Bjelajac and Filipović, 2021).
@@
 Challenges:
-Recent social media platform statements [^5] https://about.fb.com/news/2025/01/meta-more-speech-fewer-mistakes highlight the limitations of AI-based content moderation at scale.
+Recent social media platform statements [^5] https://about.fb.com/news/2025/01/meta-more-speech-fewer-mistakes highlight the limitations of AI-based content moderation.
 Challenges include the subjective nature of moderation, regional regulations, and the culturally diverse and implicit nature of content (Ocampo et al., 2023).
@@
 Opportunities:
-Opportunities for future work include building low-resource, robust, and generalizable LLM moderation frameworks that are deployable in real time, requiring stronger collaboration between researchers, moderators, and policymakers (Munzert et al., 2025; Bui et al., 2025).
-A clearer, more widely accepted taxonomy of digital violence is needed to reduce ambiguity in label definitions (Yimam et al., 2024).
-The current focus on text-based hate overlooks multimodal content like video and audio, calling for LLMs with multimodal moderation capabilities.
-To address culturally diverse and implicit content, future systems should integrate explainability, fact-checking, and real-time tools such as RAGs and search APIs.
-Finally, moderation efforts must expand beyond hate speech to cover the full spectrum of digital violence, while also embracing more inclusive approaches that consider author and receiver demographics, and broader cultural and linguistic diversity.
+We argue that future work should build low-resource, robust, and generalizable LLM moderation frameworks deployable in real time, requiring stronger collaboration between researchers, moderators, and policymakers (Munzert et al., 2025; Bui et al., 2025).
+A shared taxonomy of digital violence is needed to reduce labeling ambiguity (Yimam et al., 2024).
+Current text-centric moderation overlooks multimodal content such as video and audio, highlighting the need for multimodal LLMs.
+To handle implicit and culturally diverse cases, future systems should integrate explainability, fact-checking, and real-time tools like RAGs and search APIs.
+Finally, moderation must move beyond hate speech to encompass the broader spectrum of digital violence—including harassment, bullying, spam, and abuse—through inclusive, context-aware systems that reflect diverse cultural, regional, and linguistic realities Moghaddam et al.
+(2025); Arora et al.
+(2023).
 9 [Uncaptioned image] Misinformation
-Ranked among the top five global risks in both short- and long-term scenarios (GR5), misinformation has a direct impact on the success of many of the SDGs.
-[^6] https://www.un.org/sites/un2.un.org/files/information-integrity-and-sdgs-en.pdf NLP approaches to misinformation include tasks such as fake news detection, rumor classification, stance detection, and fact checking (Oshikawa et al., 2020; Nakashole and Mitchell, 2014).
-Early methods used stylometric features, while modern systems use neural models, pre-trained LMs, and techniques like retrieval-augmented generation and prompt-based learning in multimodal and multilingual settings (Akhtar et al., 2023; Chen and Shu, 2024b).
+Misinformation, the exposure to incorrect or misleading information, is ranked among the top five global risks in both short- and long-term scenarios (GR5), and has a direct impact on the success of many of the SDGs.
+[^6] https://www.un.org/sites/un2.un.org/files/information-integrity-and-sdgs-en.pdf NLP tools can be both part of the solution, but also part of the problem (see §10).
+There are many approaches to tackle misinformation with tasks such as fake news detection, rumor classification, stance detection, and fact checking (Oshikawa et al., 2020; Nakashole and Mitchell, 2014).
+Early methods used stylometric features, while modern systems use neural models, pre-trained LMs, and techniques like RAG and prompt-based learning in multimodal and multilingual settings (Akhtar et al., 2023; Chen and Shu, 2024b).
+Recently LLM agents are used for generating fact-checking articles (Sahnan et al., 2025), verifying complex claims (Chowdhury et al., 2025; Wang et al., 2025), and handling long-form or machine-generated texts (Xie et al., 2025; Boonsanong et al., 2025).
 Challenges:
@@
 Opportunities:
-A key opportunity lies in developing human-centered evaluation methods tailored to real-world misinformation detection tasks (Das et al., 2023a).
-This would also benefit from interdisciplinary collaboration—for example, by integrating fundamental theories from social sciences and economics (Zhou and Zafarani, 2020), or by partnering with moderators, policymakers, and professional fact-checkers (Warren et al., 2025).
-Additionally, advancing the timeliness of ground truth through LLM-agents with access to web evidence and external knowledge bases can improve adaptability to fast-evolving misinformation.
-Developing domain-agnostic features that generalize across topics, languages, modalities, and time would help track shifting deceptive styles (Chen and Shu, 2024b) and identify check-worthy claims.
+A key opportunity lies in developing human-centered evaluation methods tailored to real-world tasks (Das et al., 2023a), and strengthened through interdisciplinary collaboration–integrating social and economic theories (Zhou and Zafarani, 2020) and partnerships with moderators, policymakers, and fact-checkers (Warren et al., 2025).
+Improving the timeliness of ground truth via LLM-agents that access web evidence and external knowledge bases can enhance adaptability to fast-evolving content.
+Developing domain-agnostic features generalizable across topics, languages, and modalities can help detect shifting deceptive styles (Chen and Shu, 2024b) and identify check-worthy claims.
+The growing threat of LLM-generated misinformation remains a critical frontier (Liu et al., 2025a; Gabriel et al., 2024), underscoring the need for adaptive, accountable, and transparent detection systems.
 10 [Uncaptioned image] AI Harms
-AI harms are among the top global threats (GR5).
+AI harms, referring to negative impacts stemming from the design, deployment, or use of AI systems, are among the top global threats (GR5).
 NLP is part of the problem—but also part of the solution.
-We start with the black-box problem and follow the harm taxonomy by Weidinger et al.
-(2022) to explore NLP mitigation strategies.
-The Black Box Problem
+We first consider the black-box nature of AI systems and then follow the harm taxonomy by Weidinger et al.
+(2022) to structure our discussion and explore NLP-based mitigation strategies.
+The Black Box Problem.
 LLMs often operate as black boxes, offering limited interpretability and transparency (Hassija et al., 2024).
@@
 Visual methods like LIME (Ribeiro et al., 2016), ACE (Ghorbani et al., 2019), and heatmaps (Arras et al., 2017) highlight input relevance, while structural tools like SVCCA (Raghu et al., 2017), t-SNE, and TCAV (Kim et al., 2018) uncover concept-level insights.
-Representation and Toxicity
+Representation and Toxicity.
 NLP techniques are actively developed to address biases in training data and AI models (Gallegos et al., 2024), to reduce bias and toxicity.
 Further discussion is provided in §7 on inequality and §8 on digital toxicity.
-Privacy, Safety and Malicous Uses
-NLP approaches to harm reduction include learning to refuse or prevent memorization to protect personal data (Carlini et al., 2021; Liu et al., 2025b), adversarial training (Goyal et al., 2023), safe decoding (Xu et al., 2024b), authorship verification (Huang et al., 2024a), safety alignment (Bhardwaj et al., 2024), and red-teaming to expose vulnerabilities (Purpura et al., 2025).
-Ungrounded Knowledge
+Privacy, Safety and Malicous Uses.
+NLP approaches to harm reduction include learning to refuse or prevent memorization to protect personal data (Carlini et al., 2021; Liu et al., 2025c), adversarial training (Goyal et al., 2023), safe decoding (Xu et al., 2024b), authorship verification (Huang et al., 2024a), safety alignment (Bhardwaj et al., 2024), and red-teaming to expose vulnerabilities (Purpura et al., 2025).
+Ungrounded Knowledge.
 As pointed in §9 LLM outputs can include hallucinations and spread misinformation (Pan et al., 2023; Huang et al., 2024b; Chen and Shu, 2024b).
 Mitigation strategies include grounding with external knowledge via RAG (Lewis et al., 2020; Peng et al., 2023; Asai et al., 2024), expressing uncertainty (Yang et al., 2023c; Feng et al., 2024; Xiong et al., 2024; Deng et al., 2024), and self-verification methods (Weng et al., 2023; Hong et al., 2024).
-Socioeconomic and Environmental Harms
+Socioeconomic and Environmental Harms.
 In NLP, researchers have called for greater climate awareness and transparency through reporting frameworks (Hershcovich et al., 2022b), while others focus on reducing energy use with model optimization techniques like pruning, quantization, and distillation (Schwartz et al., 2020; Jin et al., 2024; Zhu et al., 2024).
@@
 There are also several governance problems and establishing equitable partnerships with nonprofits, where power imbalances or misaligned goals can hinder collaboration.
+Lastly, while NLP methods can mitigate certain AI harms, many of them—such as overreliance or erosion of human agency—cannot be resolved through technical solutions alone.
 Opportunities:
@@
 Multidisciplinary approaches such as logical relationships between inputs (Ayoobi et al., 2025; Freedman et al., 2025) and mechanistic interpretations of model behavior (Hou et al., 2023; Yu and Ananiadou, 2024) can further enhance transparency.
-New research directions could be implemented by studying LLM overreliance, manipulation and unfair distribution of benefits from model access.
+Additional studies could examine LLM overreliance, manipulation, the uneven distribution of benefits from model access, as well as tracing gaps and overlaps between professional and laypeople AI concerns (Karamolegkou et al., 2025a).
 Embracing process-aware NLP enables alignment with domain logic and ethical goals through explainability and fairness-by-design (Bernardi et al., 2024; Zhuang et al., 2025).
@@
 This work reviewed nine key application areas, addressing RQ1 on existing impactful solutions with concrete examples of how modern NLP technologies have already contributed.
-At the same time, in addressing RQ2 on key open challenges, we find that areas like misinformation, online harms, and education have seen sustained attention in NLP, while domains such as poverty alleviation, environmental protection, peacebuilding are only starting to gain traction in response to real-world crises.
-In response to RQ3 concerning promising directions for future work, we outline a set of universal challenges, opportunities, and actionable recommendations to guide future research in NLP4SG:
-Recurring Challenges: Across these diverse domains, we identified five systemic challenges that continue to hinder progress: (1) Data scarcity and representational bias, particularly affecting low-resource languages and marginalized groups; (2) Misaligned evaluation metrics that fail to capture human-centered qualities like empathy or cultural sensitivity; (3) Safety, privacy, and ethical concerns, which are magnified in high-stakes settings such as mental health or misinformation detection; (4) Limited generalization across cultures, modalities, and real-world conditions; and (5) Infrastructural gaps and institutional fragmentation, especially in underrepresented scenarios.
-Emerging Opportunities: Despite these barriers, we highlight several promising cross-cutting directions: (1) Multimodal and multilingual learning can help systems better reflect real-world diversity; (2) Human-AI collaboration enables more aligned, adaptive and interpretable NLP pipelines; (3) Participatory design and evaluation ensure that benchmarks and systems are co-developed with affected communities; (4) Retrieval-augmented and policy-aware methods offer new tools for building verifiable and context-sensitive applications; and (5) Explainability and AI literacy foster critical engagement and equitable access.
-Call to Action: To advance NLP4SG, we call on the community to: (1) develop joint benchmarks featuring multilingual, culturally diverse, and socially grounded data; (2) collaborate closely with domain experts, such as educators, health practitioners, and civil society organizations, to co-design evaluation frameworks that reflect end-user needs; (3) pursue human-centered methodologies instead of one-size-fits-all solutions.
+Complementing this review, our quantitative analysis of 47K ACL papers showed that research on AI harms and inclusion has grown most rapidly, while domains like poverty, peacebuilding, and environment remain underexplored (Figure 1 and Appendix A).
+At the same time, in addressing RQ2 on key open challenges, we find that areas like misinformation, online harms, and education have seen sustained attention in NLP, while domains such as poverty alleviation, environmental protection, and peacebuilding are only starting to gain traction in response to real-world crises.
+Addressing RQ3 concerning promising directions for future work, we outline a set of universal challenges, opportunities, and actionable recommendations to guide future research in NLP4SG:
+Emerging Challenges and Opportunities: Despite barriers such as data scarcity and representational bias, misaligned evaluation metrics, and persistent safety, privacy, and ethical concerns, alongside ongoing infrastructural gaps, we identify several promising directions for progress.
+(1) Multilingual and multicultural learning can help systems better reflect real-world diversity including not only mainstream languages, but caring about local language communities equally; (2) Human–AI collaboration enables more adaptive and interpretable NLP pipelines; (3) Participatory design and evaluation ensure that systems are co-developed with affected communities; (4) Retrieval-augmented and policy-aware methods provide tools for verifiable, context-sensitive applications; and (5) Explainability and AI literacy foster critical engagement and equitable access.
+Call to Action: To advance NLP4SG, we call on the community to: (1) Develop joint benchmarks featuring multilingual, culturally diverse, and socially grounded data; (2) Collaborate closely with domain experts, such as educators, health practitioners, and civil society organizations, to co-design evaluation frameworks that reflect end-user needs.
+(3) Pursue human-centered methodologies instead of one-size-fits-all solutions.
 Progress depends on pluralistic, context-aware roadmaps that align with both local realities and global development goals.
-Finally, while modern LLMs offer significant potential, it is crucial to ensure their affordability and accessibility so they serve the public good rather than exacerbate existing inequalities.
+(4) Finally, while modern LLMs offer significant potential, it is crucial to ensure their affordability and accessibility so they serve the public good rather than exacerbate existing inequalities.
+Assess whether deploying a large, one-size-fits-all generalist LLM is truly necessary, or whether more efficient, and more environmentally sustainable NLP solutions would be preferable.
 NLP has the tools to move beyond abstract benchmarking and toward socially responsive technologies designed with—and for—impacted communities.
 Realizing this vision requires not just technical innovation but also sustained interdisciplinary collaboration, inclusive practices, and a commitment to long-term global equity.
-We hope our findings can help earlier career researchers find their research niche and that more advanced researchers will have a fresh overview of the field to foster NLP4SG applications with a more interdisciplinary paradigm.
+We hope our findings can help researchers early in their careers to find their research niche and that more advanced researchers will have a fresh overview of the field to foster NLP4SG applications with a more interdisciplinary paradigm.
 Limitations
@@
 This work is grounded in the belief that NLP research should be aligned with broader societal priorities and developed with care for its downstream impact.
-In proposing mappings between NLP directions, global goals, and risks, we are mindful of the potential for unintended consequences and overgeneralization.
+All ACL materials used for our statistical analysis are licensed using the Creative Commons 3.0 BY-NC-SA.
+In proposing mappings between NLP directions, global goals, and risks, we are mindful of the potential for unintended consequences.
 The open areas of work mentioned throughout the paper are just suggestive, and the practitioners, when working on them, should carefully evaluate the setting and downstream impacts.
@@
 Where possible, we cite and build on existing initiatives in NLP for social good, and we support efforts to foreground equity, inclusivity, and transparency in both research framing and methodology.
+Throughout this paper, we emphasize that NLP systems should augment rather than replace human expertise, particularly in sensitive or high-stakes contexts.
 Finally, we acknowledge that AI assistants were used for proofreading this paper, a practice we consider ethically acceptable under appropriate usage.
@@
 Zhijing Jin was supported partly by the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039B; by the Machine Learning Cluster of Excellence, EXC number 2064/1 – Project number 390727645; and by Schmidt Sciences SAFE-AI Grant.
-Daryna Dementieva’s work was additionally supported by the Friedrich Schiedel TUM Fellowship.
-Appendix A Paper Selection: Thematic Tables of Datasets, NLP Tasks, Evaluation Metrics, and References
-A.1 Health and Well-being
+Daryna Dementieva’s and Alexander Fraser’s work was supported by the European Research Council (ERC) under grant agreement No.
+101113091 - Data4ML, an ERC Proof of Concept Grant as well as co-funded by the an ERC Advanced Grant EPICAL, grant agreement No.
+101141712.
+Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council.
+Neither the European Union nor the granting authority can be held responsible for them.
+Daryna Dementieva’s work was additionally supported by the Friedrich Schiedel TUM Think Tank Fellowship.
+Appendix A Quantitative Analysis
+A.1 Data Preparation
+The PaperAnalyzer framework (Adauto et al., 2023) introduced a large-scale methodology for classifying research papers into socially relevant categories, drawing connections between NLP tasks, methods, and broader social goals such as the United Nations Sustainable Development Goals (SDGs).
+By leveraging large language models (LLMs) for annotation, PaperAnalyzer demonstrated how automated approaches can reduce the dependence on costly manual labeling while enabling scalable analysis of research trends.
+Our work is inspired by this methodology and adapts it to a more focused investigation of yearly trends within the ACL Anthology.
+Specifically, we designed an annotation pipeline to assign papers to the set of key NLP application areas, enabling us to capture how the community’s research priorities have evolved over time.
+For this study, we focused on ACL Anthology papers published between 2019 and 2024, including both main conference and workshop proceedings.
+This filtering yielded 47,078 papers out of the 113,207 entries in the Anthology.
+Across this subset, approximately 7% of papers lack abstracts; these papers were retained, with only their titles used for annotation.
+A.2 Annotation of Key NLP Domains
+A.2.1 Methodology
+To analyze the research landscape from a social good perspective, we use a zero-shot annotation pipeline that categorized papers into nine NLP application areas: Healthcare, Education, Poverty, Peacebuilding, Environment Protection, Inclusion & Inequalities, Online Harms, Misinformation, and AI Harms.
+Papers could be assigned to multiple domains when appropriate, while those not fitting any category were labeled as Unrelated.
+Each paper was evaluated using its title and abstract (when available).
+The model was explicitly instructed with the full list of domains, together with concise descriptions for each.
+These descriptions were initially drafted with the assistance of ChatGPT, but were subsequently reviewed and edited by the authors to ensure contextual appropriateness for the study.
+Our prompt can be viewed in Figure 5.
+For the large-scale annotation of the full dataset, we used OpenAI’s GPT-4.1 mini (openai/gpt-4.1-mini-2025-04-14), which provided both efficiency and value at scale.
+All queries were executed with temperature = 0 to ensure deterministic and reproducible outputs.
+The pipeline was applied to the complete set of 47,078 ACL Anthology papers published between 2019 and 2024.
+The total cost to annotate all papers was approximately \$12 USD.
+A.2.2 Results
+Overall, the pipeline annotated 36,030 papers (approximately 76.5%) with at least one valid domain label, with each of these papers receiving an average of 1.82 domain assignments.
+During annotation, 109 papers received at least one invalid annotation.
+Of these, 107 papers contained a mix of valid and invalid annotations, which we handled by retaining only the valid domain labels.
+The remaining 2 papers contained exclusively invalid annotations and were therefore treated as Unrelated.
+To examine how social good research is distributed across different publication venues, we analyzed the distribution of the key NLP domains across conference papers and workshop papers separately, as shown in Figure 4.
+Conferences were identified by filtering for entries where “Conference,” “Findings,” or “Annual Meeting” appeared in the booktitle field, yielding 32,886 papers.
+Workshop papers were identified by filtering for entries where “Workshop” appeared in the booktitle field, yielding 12,723 papers.
+Figure 3: Co-occurrence heatmap of annotated NLP domains.
+(a) Conference Papers
+(b) Workshop Papers
+Figure 4: Overview of the normalized NLP domain distributions for conference papers ($n=32,886$) and workshop papers ($n=12,723$) from 2019 to 2024.
+Based on these graphs, workshops show more pronounced year-to-year fluctuations and a stronger emphasis on emerging or experimental areas.
+While AI Harms and Inclusion & Inequalities remain dominant in both, workshop papers show relatively higher proportions of Online Harms and Misinformation, suggesting that exploratory and early-stage research on socially sensitive topics tends to appear more frequently in workshop venues before reaching conference-level maturity.
+Overall, the co-occurrence heatmap in Figure 3 confirms these broader patterns, showing that research on AI Harms and Inclusion & Inequalities often overlaps with multiple domains, making them central themes in NLP4SG.
+By contrast, areas such as Poverty and Peacebuilding appear more isolated, suggesting that while ethical and fairness-related work has gained strong interdisciplinary traction, other social domains remain comparatively underexplored.
+LLM Annotation Prompt You will be given the title and abstract of a research paper.
+Your task is to annotate the paper with the relevant social impact domains.
+- Review the list of domains provided.
+- Select all domains that are relevant to the paper based on its title and abstract; you can select multiple.
+- If none of the domains apply, respond with "Unrelated".
+- Do not explain your choices.
+- Do not include anything other than the list of exact domain names or the word "Unrelated".
+**Domains**: Healthcare – Enhance physical and mental well-being by improving access, diagnosis, and support through language technologies.
+Education – Ensure inclusive and equitable quality education with NLP tools for learning, teaching, and accessibility.
+Poverty – Reduce economic hardship by enabling data-driven policies, resource distribution, and financial inclusion with NLP.
+Peacebuilding – Promote peaceful societies by using NLP for dynamic conflict analysis, human rights protection, and crisis response through early-warning and mediation support.
+Environment Protection – Address environmental risks by monitoring and analyzing them, and supporting sustainability, climate action and environmental responsibility using NLP techniques.
+Inclusion and Inequalities – Foster fairness and representation by using NLP to address bias, improve accessibility, and empower marginalized voices.
+Online Harms – Make digital spaces safe by detecting and mitigating hate speech, harassment, and harmful online behaviors.
+Misinformation – Safeguard information integrity by detecting, verifying, and countering false and misleading content.
+AI Harms – Build responsible AI by improving transparency, accountability, and fairness to prevent unintended harm from NLP systems.
+**Paper Title**: {paper_title} **Paper Abstract**: {paper_abstract} Please respond in the following format only: Domain; Domain; … OR Unrelated
+Figure 5: Prompt used to annotate ACL Anthology papers with the key NLP research directions.
+A.3 Annotation of NLP Tasks and Methods
+Building on the approach proposed in Task 3 of PaperAnalyzer (Adauto et al., 2023), which analyzed NLP research for social good through its underlying tasks and methods, we aimed to reproduce a similar insight.
+The goal was to capture how specific techniques and problem types evolve across socially relevant research areas, complementing our domain-level annotation.
+A.3.1 Methodology
+We began with the set of tasks and methods previously generated from our 900-paper sample subset.
+We prompted Gemma 3 (google/gemma-3-27b-it) to identify and extract the tasks and methods mentioned in each paper’s title and abstract.
+This process yielded approximately 750 unfiltered task terms and nearly 1,000 unfiltered method terms.
+To consolidate these into a more structured taxonomy, we provided all extracted terms to ChatGPT and asked it to group them into canonical categories representing the most common NLP tasks and methods.
+We then manually reviewed each category to ensure correctness.
+The resulting taxonomy included 45 distinct task categories and 49 method categories.
+These categories were then used to annotate the full set dataset, where we used Google’s Gemini 2.5 Flash (google/gemini-2.5-flash) to assign one or more task and method labels to each paper based on its title and abstract.
+The prompt used for this extraction (shown in Figure 8) instructed the model to return only relevant tasks and methods in a standardized format, ensuring consistency across the dataset.
+The total cost to annotate all papers was approximately \$22 USD.
+A.3.2 Results
+After the annotation process, only one paper contained an invalid task label, which was manually removed.
+Sixteen papers contained no valid method annotations; for these, the method field was replaced with the label Unknown.
+Overall, task coverage was nearly complete, with 46,996 papers (99.83%) successfully assigned one or more task labels and an average of 2.74 tasks per paper.
+Method coverage was similarly high, with 45,034 papers (95.66%) receiving at least one method label and an average of 2.22 methods per paper.
+Finally, to visualize the relationships between the annotated categories, we generated a Sankey diagram (Figure 6) linking the domains, tasks, and methods, illustrating how research themes and techniques intersect across the ACL Anthology corpus.
+Figure 6: Diagram showing the flow of research focus from the key NLP domains to the top 15 NLP tasks and the corresponding top 10 methodological approaches.
+The diagram reveals that recent NLP4SG research has increasingly centered on AI Harms and Inclusion & Inequalities, highlighting a shift toward fairness, bias mitigation, and ethical analysis of large language models.
+These domains are most frequently associated with Dataset Creation, Model Analysis & Interpretability and Text Classification, indicating a growing focus on building representative data, evaluating model behavior and developing practical applications for social impact.
+Together, the task and method trends (Figure 7) illustrate a clear evolution in NLP4SG research.
+Tasks such as Model Analysis & Interpretability, Dataset Creation & Annotation, and Text Classification have gained prominence, reflecting a move toward model evaluation and analysis.
+Meanwhile, Transformers and Transfer Learning are gradually giving way to approaches like Prompting and In-Context Learning, highlighting the growing integration of LLM-driven, instruction-based methodologies.
+(a) Tasks
+(b) Methods
+Figure 7: Overview of the normalized NLP task and method distributions for ACL Anthology papers from 2019 to 2024.
+Each plot shows trendlines for the 14 most frequent tasks and methods, selected based on their average frequency across all years.
+LLM Annotation Prompt You will be given the title and abstract of a research paper.
+Your task is to annotate the paper with the primary NLP task(s) and method(s) that this paper is dealing with.
+- Review the list of tasks and methods provided.
+- Select all tasks and methods that are relevant to the paper based on its title and abstract; you can select multiple.
+- If none of the tasks or methods apply, respond with "Unknown".
+- Do not explain your choices.
+- Do not include anything other than the list of task or method labels or the word "Unknown".
+- Make sure to respond with the exact label for the tasks and methods.
+Do not modify them in any way.
+**Task List**: T1.
+Text Classification T2.
+Sequence Labeling T44.
+Web Interaction & Automation T45.
+Benchmarking & Comparative Analysis **Method List**: M1.
+Recurrent Neural Networks (RNNs, LSTMs, GRUs) M2.
+Transformers & Attention Mechanisms M48.
+Classical ML (Logistic Regression, SVM, CRF) M49.
+Tool Integration & API Augmentation **Paper Title**: {paper_title} **Paper Abstract**: {paper_abstract} Please respond in the following format **only**: Tasks: T#; T#; … Methods: M#; M#; …
+Figure 8: Prompt used to annotate ACL Anthology papers with NLP tasks and methods.
+Appendix B Paper Selection: Thematic Tables of Datasets, NLP Tasks, Evaluation Metrics, and References
+B.1 Health and Well-being
 Mental health is a key component of human health, encompassing our emotional, psychological, and social well-being SAMHSA (2023).
@@
 (2024); Liu et al.
-(2025a); Hsu et al.
+(2025b); Hsu et al.
 (2025); Kampman et al.
@@
 Automatic evaluation on symptom severity, cognitive distortion, and overall depression Severity Liu et al.
-(2025a)
+(2025b)
 MIDAS Publicly available dataset in mental health counseling Expert evaluation, reflection to question ratio, accuracy, F-1 Gunal et al. (2025)
@@
 ELMTEX Dataset (clinical summaries) Information extraction, Entity normalization ROGUE, BERTScore, Precision, Recall, F1 Guluzade et al. (2025)
-A.2 Education
+B.2 Education
 NLP for Education Tools & Systems.
@@
 National Center for Teacher Effectiveness (NCTE) Text Classification (Evaluation of Classifying Educational Discourse Features) Accuracy, Precision, Recall, F1 Demszky and Hill (2023)
-A.2.1 AI Literacy
+B.2.1 AI Literacy
 To identify relevant literature on AI literacy, we used Semantic Scholar and Google Scholar search terms such as “ai literacy” and (“ai literacy” + “social impact” + “nlp”).
 We found most of the papers selected for this topic to be on classroom-based studies and measurements and metrics for AI literacy, along with some interdisciplinary papers connecting AI literacy to other disciplines, for example, psychology.
-A.3 Peace Building
+B.3 Peace Building
 We have organized highlighted papers found in our review of papers at the intersection of peace building and NLP in Table 3.
@@
 Table 3: Overview of Peace Building and Physical Safety related NLP studies
-A.4 Poverty
+B.4 Poverty
 We began our review of papers focused on the study of poverty by through keyword searches such as “poverty detection" on Google Scholar.
@@
 Table 4: Overview of Poverty related NLP studies
-A.5 Online Harms
+B.5 Online Harms
 Initial research primarily focused on developing classification frameworks across different spectrum of online harms based on existing AI models.
@@
 Table 5: Representative datasets, NLP task(s) and evaluation metrics from research on online harm.
-A.6 Misinformation
+B.6 Misinformation
 Methodology.
@@
 Table 6: Datasets on misinformation detection.
-A.7 Inequalities and Bias
+B.7 Inequalities and Bias
 Datasets NLP Task(s) Key Evaluation Metrics Reference
@@
 BiosBias (LinkedIn biographies) Occupation classification F1 diff., error disparity De-Arteaga et al. (2019)
-LLM-Agent Interaction Logs Multi-agent bias propagation Bias-Score, fairness gap Borah and Mihalcea (2024)
-BiosBias+GPT-J Causal weight-editing for debiasing Stereotype score, accuracy Cai et al. (2024)
-TED-Talk / IWSLTEn–It MT Pronoun-drop attribution maps Pronoun-drop rate, TER gap Attanasio et al. (2023)
-Synthetic counterfactual pairs (Iter CDA) Bias-robust text classification Bias amplification ratio, F1 Plyler and Chi (2025)
-GeoWAC, Reddit, UN General Debates Region-aware bias evaluation metric Region-aware WEAT effect size, mismatchBorah et al. (2025)
-CrowS-Pairs (intersectional ext.) Intrinsic bias evaluation Bias Score (Black vs White) Guo and Caliskan (2021)
+LLM‑Agent Interaction Logs Multi‑agent bias propagation Bias‑Score, fairness gap Borah and Mihalcea (2024)
+BiosBias+GPT‑J Causal weight‑editing for debiasing Stereotype score, accuracy Cai et al. (2024)
+TED‑Talk / IWSLTEn–It MT Pronoun‑drop attribution maps Pronoun‑drop rate, TER gap Attanasio et al. (2023)
+Synthetic counterfactual pairs (Iter CDA) Bias‑robust text classification Bias amplification ratio, F1 Plyler and Chi (2025)
+GeoWAC, Reddit, UN General Debates Region-aware bias evaluation metric Region‑aware WEAT effect size, mismatch% Borah et al. (2025)
+CrowS‑Pairs (intersectional ext.) Intrinsic bias evaluation Bias Score (Black vs White) Guo and Caliskan (2021)
 Table 7: Representative datasets and evaluation practices across gender-bias NLP tasks.
@@
 FORK test set Commonsense QA (Culinary) Accuracy on US vs non-US probes; Statistical significance Palta and Rudinger (2023)
-TV show dialogues; Cross-culture shows; LDC conversational corpora Norm Extraction; Self-Verification; Grounding AUC for grounding; human-judged best-norm selection Fung et al. (2024)
+TV show dialogues; Cross‑culture shows; LDC conversational corpora Norm Extraction; Self-Verification; Grounding AUC for grounding; human-judged best-norm selection Fung et al. (2024)
 Reddit corpus of 61,981 users; Word association benchmarks Demographic-Conditioned LM; Word Association Perplexity; word-association accuracy Welch et al. (2020)
@@
 Table 9: Representative datasets and evaluation practices across accessibility-related NLP tasks.
-A.8 Environmental Harms
-Given the absence of a survey paper in this field, we began our review with research presented at the inaugural https://aclanthology.org/events/climatenlp-2024/.
+B.8 Environmental Harms
+Given the absence of a survey paper in this field, we began our review with research presented at the inaugural Workshop on Natural Language Processing Meets Climate Change (ClimateNLP 2024).
 Additionally, we have included follow-up studies conducted by researchers within the ClimateNLP community.
@@
 Table 10: Selection of influential datasets and papers in the ClimateNLP domain, ranging from topic modelling over text classification to question answering and information extraction
-Appendix B Global Goals
+Appendix C Global Goals
 We include an overview of the Sustainable Development Goals (SDGs) in Figure 9, which apply to many of the NLP applications discussed in this paper.
 Figure 9: Overview of the SDG goals. Source: https://sdgs.un.org/goals
-Appendix C Global Risks
+Appendix D Global Risks
 We present the key global risks categorized by domain, as outlined in the Global Risks Report 2025 by the World Economic Forum.
@@
 Figure 11: Global risks ranked by severity over the short and long term. Reproduction of Figure FIGURE C from: the Global Risks Report 2025
-Appendix D Guidelines for authors
+Appendix E Guidelines for authors
 This work involved contributions from many collaborators with diverse backgrounds, who surveyed various topics and guided the writing process.